### v1

# Alternating Gradient-Type Algorithm for Bilevel Optimization with Inexact Lower-Level Solutions via Moreau Envelope-based Reformulation

Abstract
In this paper, we study a class of bilevel optimization problems where the lower-level problem is a convex composite optimization model, which arises in various applications, including bilevel hyperparameter selection for regularized regression models. To solve these problems, we propose an Alternating Gradient-type algorithm with Inexact Lower-level Solutions (AGILS) based on a Moreau envelope-based reformulation of the bilevel optimization problem. The proposed algorithm does not require exact solutions of the lower-level problem at each iteration, improving computational efficiency. We prove the convergence of AGILS to stationary points and, under the Kurdyka-Łojasiewicz (KL) property, establish its sequential convergence. Numerical experiments, including a toy example and a bilevel hyperparameter selection problem for the sparse group Lasso model, demonstrate the effectiveness of the proposed AGILS.
Keywords:
Bilevel optimization, hyperparameter selection, inexact, alternating gradient descent, convergence, sparse group Lasso
1 Introduction
Bilevel optimization problems are a class of hierarchical optimization problems where the feasible region is implicitly defined by the solution set of a lower-level optimization problem. In this work, we focus on a bilevel optimization problem which has a convex composite lower-level problem,
$\min\limits_{x\in X,y\in Y}$ $\ F(x,y)$ (1)
${\rm s.t.}$ $\ y\in S(x),$
where $S(x)$ denotes the set of optimal solutions to the lower-level problem:
$\min\limits_{y\in Y}\ \varphi(x,y):=f(x,y)+g(x,y).$ (2)
Here, $X\in{\mathbb{R}}^{n}$ and $Y\in{\mathbb{R}}^{m}$ are closed convex sets, $f(x,y):{\mathbb{R}}^{n}\times{\mathbb{R}}^{m}\rightarrow{\mathbb{R}}$ is smooth and convex with respect to $y$ on $Y$ for any $x\in X$, and $g(x,y):{\mathbb{R}}^{n}\times{\mathbb{R}}^{m}\rightarrow{\mathbb{R}}$ is convex but potentially nonsmooth with respect to $y$ on $Y$ for any $x\in X$. Additionally, we assume that $g(x,\cdot):{\mathbb{R}}^{m}\rightarrow{\mathbb{R}}$ is proximal-friendly, enabling efficient evaluation of its proximal operator.
Bilevel optimization originates from economic game theory, particularly the Stackelberg game [1], and has since found diverse applications in economics [2, 3, 4], transportation [5, 6], and machine learning [7, 8, 9, 10, 11]. Notably, many hyperparameter optimization problems can be modeled as bilevel optimization problems with a convex composite lower-level problem as in (2). Bilevel optimization has been extensively studied, with monographs [12, 13, 14, 15] providing a comprehensive overview of its methodologies and applications. For bilevel problems where all defining functions are smooth, the first-order optimality condition is used to reformulate the problem as a mathematical program with equilibrium constraints (MPEC); see, e.g., [16, 17, 18, 19]. However, this approach often encounters difficulties when applied to larger-scale problems or those involving non-smooth lower-level functions, such as those with regularization terms. Smoothing techniques have also been proposed to address nonsmooth issues in specific applications, such as hyperparameter optimization [20, 21]. Another widely used approach for reformulating bilevel optimization problems is based on the value function [22, 23]. Specifically, the bilevel problem is reformulated as the following equivalent problem,
$({\rm VP})\quad\quad\min\limits_{x\in X,y\in Y}\ F(x,y)\quad{\rm s.t.}\quad\ \varphi(x,y)-v(x)\leq 0,$ (3)
where $v(x):=\inf_{\theta\in Y}\varphi(x,\theta)$ is the value function of the lower-level problem. Under partial calmness condition, Newton-type methods have been developed for this reformulation [24, 25, 26]. For fully convex lower-level problems, (VP) can be treated as a difference-of-convex (DC) program, as studied in [27], which led to applications in bilevel hyperparameter tuning problems [28]. Subsequently, [29] extended these ideas to settings where the fully convex assumption is relaxed, introducing a Moreau envelope-based reformulation and corresponding algorithms.
In recent years, bilevel optimization has garnered significant attention within the machine learning community, driving the development of various methods and algorithms. Among these, gradient-based approaches have emerged as particularly effective due to their simplicity and efficiency [8, 10, 11]. Many of these methods are built upon the computation of hypergradient or its approximation, especially in cases where the lower-level problem is smooth and strongly convex. The strong convexity of the lower-level problem guarantees the uniqueness of its solution, facilitating hypergradient computation [30, 31, 32, 33, 34, 35]. Among these, [34] proposed a two-timescale single-loop stochastic gradient algorithm (TTSA) that efficiently tackles smooth bilevel problems. In addition, recent works have proposed gradient-based approaches based on the value function reformulation (VP) of bilevel problems, see, e.g., [36, 37, 38, 39, 40]. These approaches typically require smoothness conditions for the lower-level problem. Notably, [39] firstly proposed first-order penalty methods for finding $\epsilon$-KKT solutions of (VP) via solving a minimax optimization problem.
A key challenge in gradient-type algorithms derived from the value function reformulation is the computation of the gradient or subgradient of the value function at each iteration [28, 27]. This involves solving the corresponding lower-level problem, even when the gradient exists, as in the case where $\nabla v(x)=\nabla_{x}\varphi(x,y^{*}(x))$, with $y^{*}(x)=\arg\min_{\theta\in Y}\varphi(x,\theta)$ ([41, Remark 4.14]). However, exactly solving the lower-level problem at each iteration can be computationally expensive, potentially limiting the efficiency of the algorithm. In cases where the lower-level problem satisfies uniformly strong convexity or a global PL condition, recent works [37, 38] use inexact solutions to the lower-level problem, replacing the exact solution $y^{*}(x)$ in the gradient computation, constructing an approximation to the value function gradient. But, when the lower-level problem lacks uniform strong convexity and a global PL condition, constructing an approximation to $\nabla v$ using inexact lower-level solutions becomes challenging. Specifically, using an inexact lower-level solution $y_{k}$ with arbitrarily small first-order residual (i.e., $\|\nabla_{y}f(x_{k},y_{k})\|\rightarrow 0$) still results in a fixed gap between the approximate gradient $\nabla_{x}\varphi(x_{k},y_{k})$ and the true gradient $\nabla v(x_{k})$, even as $k\rightarrow\infty$. As an illustrative example, consider the following simple lower-level problem,
$\min_{y\in\mathbb{R}}\varphi(x,y):=xy^{2},$ (4)
where $x,y\in\mathbb{R}$. For $x>0$, $y^{*}(x)=0$, so $v(x)=0$ and $\nabla v(x)=0$. However, for any vanishing sequence $\epsilon_{k}\rightarrow 0$, consider sequences $x_{k}={\epsilon_{k}}/{2}$ and $y_{k}=1$. Here, $y_{k}$ is an inexact solution to (4) for $x=x_{k}$, with a small first-order residual $\|\nabla_{y}\varphi(x_{k},y_{k})\|=\epsilon_{k}$, which can be arbitrarily small as $k\rightarrow\infty$. Despite this, the gradient approximation $\nabla_{x}\varphi(x_{k},y_{k})$ will still have a constant gap to the true gradient $\nabla v(x_{k})$, since $\|\nabla_{x}\varphi(x_{k},y_{k})-\nabla v(x_{k})\|=1$ even as $k\rightarrow\infty$. This example demonstrates the difficulty of constructing accurate value function gradient approximations and developing gradient-type algorithms with inexact solutions to the lower-level problem for bilevel optimization when the lower-level problem does not have uniformly strong convexity.
To overcome this challenge, in this paper, we consider the following Moreau envelope-based reformulation of the bilevel optimization problem (1), first introduced in [29],
$({\rm VP})_{\gamma}\quad\quad\min\limits_{x\in X,y\in Y}\ F(x,y)\quad{\rm s.t.}\quad\ \varphi(x,y)-v_{\gamma}(x,y)\leq 0,$ (5)
where $v_{\gamma}(x,y)$ is the Moreau envelope associated with the lower-level problem,
$v_{\gamma}(x,y):=\inf\limits_{\theta\in Y}\varphi(x,\theta)+\frac{1}{2\gamma}\|\theta-y\|^{2},$ (6)
where $\gamma>0$. When the lower-level problem is convex with respect to $y$, this reformulation is equivalent to the original bilevel optimization problem (1) (see [29, Theorem 1]). Directional optimality conditions for $({\rm VP})_{\gamma}$ are discussed in [42]. However, as shown in [29] and [27, Proposition 7], classical constraint qualification (CQ) conditions do not hold at any feasible point of $({\rm VP})_{\gamma}$, motivating us to study the following relaxed approximation problem,
$({\rm VP})_{\gamma}^{\epsilon}\quad\quad\min\limits_{x\in X,y\in Y}\ F(x,y)\quad{\rm s.t.}\quad\ \varphi(x,y)-v_{\gamma}(x,y)\leq\epsilon,$ (7)
where $\epsilon>0$. [29, Proposition 6] shows that for any $\epsilon_{0}>0$, there exists $\epsilon>0$ such that a local minimizer of $({\rm VP})_{\gamma}^{\epsilon}$ is $\epsilon_{0}$-close to the solution set of the original bilevel optimization problem (1). In [29], a double-loop difference-of-convex algorithm was developed to solve $({\rm VP})_{\gamma}^{\epsilon}$ for general bilevel optimization problems. This algorithm requires solving two optimization subproblems at each iteration, including the proximal lower-level problem (6). Several subsequent works [43, 44, 45] have adopted ideas akin to the Moreau envelope-based reformulation to develop single-loop gradient-based bilevel algorithms. However, these works limit their convergence analysis to the decreasing property of certain merit functions, lacking rigorous theoretical support for the convergence of iterates.
In this paper, we propose an Alternating Gradient-type algorithm with Inexact Lower-level Solutions (AGILS), for solving the approximation bilevel problem $({\rm VP})_{\gamma}^{\epsilon}$. The main contributions of this paper are summarized as follows.
  • We propose AGILS, which is an alternating gradient-type algorithm with an adaptive penalty parameter update strategy for solving the approximation bilevel optimization problem $({\rm VP})_{\gamma}^{\epsilon}$. A key feature of AGILS is the introduction of a verifiable inexact criterion for the proximal lower-level problem (6). Unlike the double-loop difference-of-convex algorithm developed in [29], which requires exact solution to the proximal lower-level problem at each iteration, AGILS allows for inexact solutions, enhancing both the flexibility and computational efficiency of the algorithm. In particular, the inexact criterion also facilitates the integration of various efficient methods for solving convex composite models within the proximal lower-level problem.
  • We establish convergence results for the proposed AGILS under mild assumptions, along with clearly defined and estimable step size ranges. Specifically, we show that AGILS subsequentially converges to a Karush-Kuhn-Tucker (KKT) stationary point of $({\rm VP})_{\gamma}^{\epsilon}$. Moreover, under the Kurdyka-Łojasiewicz (KL) property, we establish the sequential convergence of the algorithm. This analysis is non-trivial due to the inexactness in solving the proximal lower-level problem, the alternating update scheme, and the lack of Lipschitz continuity of $\nabla v_{\gamma}$. Through careful analysis and the introduction of a new merit function, we successfully establish sequential convergence.
  • The proposed AGILS can efficiently handle high-dimensional bilevel optimization problems, benefiting from its use of only an inexact solution to the proximal lower-level problem (6) and an alternating gradient-type update scheme. To evaluate its performance, we conduct numerical experiments on a toy example and the sparse group Lasso bilevel hyperparameter selection problem. The numerical results demonstrate the efficiency of the AGILS compared to other commonly used approaches.
This paper is organized as follows. In Section 2, we provide the basic assumptions, and properties of the Moreau envelope-based reformulation. Section 3 presents the proposed AGILS for solving the approximation bilevel optimization problem $({\rm VP})_{\gamma}^{\epsilon}$. Section 4 provides the convergence analysis towards KKT stationary points, and Section 5 establishes the sequential convergence under the KL property. Finally, Section 6 presents numerical experiments on a toy example and the sparse group Lasso bilevel hyperparameter selection problem.
2 Preliminaries
2.1 Notations
[…]
^$\hat{\partial}h(\bar{z})$ =$=\left\{v\in\mathbb{R}^{n}\;\middle|\;h(z)\geq h(\bar{z})+\langle v,z-\bar{z}\rangle+o(\|z-\bar{z}\|),\;\forall z\in\mathbb{R}^{n}\right\},$
[…]
2.2 Basic Assumptions
This part outlines the assumptions on the problem data for the bilevel optimization problem studied in this work. First, we assume that the upper-level objective $F$ is smooth and bounded below.
Assumption 1.
The function $F:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ is bounded below on $X\times Y$, i.e., $\underline{F}:=\inf_{(x,y)\in X\times Y}F(x,y)>-\infty$. $F$ is continuously differentiable, and its gradients $\nabla_{x}F$ and $\nabla_{y}F$ are $L_{F_{x}}$- and $L_{F_{y}}$-Lipschitz continuous on $X\times Y$, respectively.
Throughout this paper, we focus on cases where the lower-level problem is a convex optimization problem.
Assumption 2.
For any $x\in X$, the functions $f(x,\cdot)$ and $g(x,\cdot)$ are convex and defined on the closed convex set $Y$.
Smoothness and weak convexity conditions are required for the lower-level objective functions. We recall that a function $h:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ is $(\rho_{1},\rho_{2})$-weakly convex on a convex set $X\times Y$ if $h(x,y)+\frac{\rho_{1}}{2}\|x\|^{2}+\frac{\rho_{2}}{2}\|y\|^{2}$ is convex on $X\times Y$. This notion generalizes convexity ($\rho_{1}=\rho_{2}=0$). Weakly convex functions [46], extend the class of convex functions to include many nonconvex functions. For weakly convex functions, the Fréchet (regular) subdifferential coincides with both the limiting subdifferential and the Clarke subdifferential (see, e.g., [47, Proposition 3.1 and Theorem 3.6]).
Assumption 3.
The lower-level objective functions $f(x,y)$ and $g(x,y)$ satisfy:
  • (1)
    $f(x,y)$ is continuously differentiable on $X\times Y$, and its gradients $\nabla_{x}f$ and $\nabla_{y}f$ are $L_{f_{x}}$- and $L_{f_{y}}$-Lipschitz continuous on $X\times Y$, respectively;
  • (2)
    $g(x,y)$ is continuous on $X\times Y$, $\nabla_{x}g(x,y)$ exists, and $\nabla_{x}g(\cdot,y)$ is $L_{g_{1}}$-Lipschitz continuous on $X$ for any $y\in Y$, while $\nabla_{x}g(x,\cdot)$ is $L_{g_{2}}$-Lipschitz continuous on $Y$ for any $x\in X$;
  • (3)
    $f(x,y)$ is $(\rho_{f_{1}},\rho_{f_{2}})$-weakly convex on $X\times Y$, and $g(x,y)$ is $(\rho_{g_{1}},\rho_{g_{2}})$-weakly convex on $X\times Y$.
Remark 1.
The bilevel problem (1) has also been studied in [43]. However, in addition to Assumption 3, Assumption 3.2 in [43] imposes an additional Lipschitz-like continuity condition on the proximal operator $\operatorname{Prox}_{{g}(x,\cdot)}$ with respect to $x$. This assumption is challenging to verify in practice and is not required in this work.
Assumption 3(1) implies $(\rho_{f_{1}},\rho_{f_{2}})$-weak convexity with $\rho_{f_{1}}=\rho_{f_{2}}=L_{f}$ of $f(x,y)$ [48, Lemma 5.7]. Furthermore, the lower-level objective $\varphi(x,y)$ is $(\rho_{\varphi_{1}},\rho_{\varphi_{2}})$-weakly convex on $X\times Y$, where $\rho_{\varphi_{1}}=\rho_{f_{1}}+\rho_{g_{1}}$ and $\rho_{\varphi_{2}}=\rho_{f_{2}}+\rho_{g_{2}}$.
According to [49, Lemma 4.2], $g(x,y)=h(c(x,y))$ is $L_{h}L_{c}$-weakly convex if $h:\mathbb{R}^{d}\to\mathbb{R}$ is a closed convex $L_{h}$-Lipschitz continuous function and $c:\mathbb{R}^{n+m}\to\mathbb{R}^{d}$ is a smooth function with an $L_{c}$-Lipschitz continuous Jacobian. Below, we provide more specific examples of $g(x,y)$ satisfying 3.
Lemma 1.
Let $g(x,y)$ satisfy one of the following conditions. Then $g(x,y)$ satisfies Assumption 3.
  • (1)
    $g(x,y)=x\|y\|_{1}$. Then $g(x,y)$ is $(m,1)$-weakly convex on +×$\mathbb{R}_{+}\times\mathbb{R}^{m}$.
  • (2)
    $g(x,y)=\sum_{j=1}^{J}x_{j}\|y^{(j)}\|_{2}$, where $y^{(j)}$ denotes the $j$-th group of $y$. Then $g(x,y)$ is $(1,1)$-weakly convex on +$\mathbb{R}^{J}_{+}\times\mathbb{R}^{m}$.
  • (3)
    $g(x,y)=xP(y)$, where $P(y)$ is convex and $L_{P}$-Lipschitz continuous on $Y$. Then $g(x,y)$ is $(\rho_{g_{1}},\rho_{g_{2}})$-weakly convex on +×$\mathbb{R}_{+}\times\mathbb{R}^{m}$ for any $\rho_{g_{1}},\rho_{g_{2}}\geq 0$ satisfying $\rho_{g_{1}}\rho_{g_{2}}\geq L_{P}^{2}$.
Proof.
For (1), the proof follows a similar discussion as in Section 6.1 of [29], using a different estimation of weak convexity constants. We note that $x\|y\|_{1}+\frac{m}{2}x^{2}+\frac{1}{2}\|y\|^{2}_{2}=\sum_{i=1}^{m}\frac{1}{2}\left(x+|y_{i}|\right)^{2}$, which is the composition of a linear function with a convex function. By [50, Proposition 1.54], the sum is convex, implying $x\|y\|_{1}$ is $(m,1)$-weakly convex. For (2), see Section 6.2 of [29]. To establish (3), we recall that $g(x,y)$ is $(\rho_{g_{1}},\rho_{g_{2}})$-weakly convex on $X\times Y$ if and only if, for any points $(x_{1},y_{1}),(x_{2},y_{2})\in X\times Y$ and $a\in[0,1]$, the approximate secant inequality holds: $g(ax_{1}+(1-a)x_{2},ay_{1}+(1-a)y_{2})\leq ag(x_{1},y_{1})+(1-a)g(x_{2},y_{2})+a(1-a)/2\cdot(\rho_{g_{1}}\|x_{1}-x_{2}\|^{2}+\rho_{g_{2}}\|y_{1}-y_{2}\|^{2})$. Now, consider $g(x,y)=xP(y)$. For any points $(x_{1},y_{1}),(x_{2},y_{2})\in\mathbb{R}_{+}\times\mathbb{R}^{m}$ and $a\in[0,1]$, we have
$g(ax_{1}+(1-a)x_{2},ay_{1}+(1-a)y_{2})-ag(x_{1},y_{1})+(1-a)g(x_{2},y_{2})$
$=$ $(ax_{1}+(1-a)x_{2})P(ay_{1}+(1-a)y_{2})-ax_{1}P(y_{1})-(1-a)x_{2}P(y_{2})$
$\leq$ $a(a-1)x_{1}P(y_{1})+a(1-a)x_{1}P(y_{2})+a(1-a)x_{2}P(y_{1})+a(a-1)x_{2}P(y_{2})$
$=$ $-a(1-a)\left(x_{1}-x_{2}\right)\left(P(y_{1})-P(y_{2})\right)$
$\leq$ $a(1-a)L_{P}\|x_{1}-x_{2}\|\left\|y_{1}-y_{2}\right\|.$
where the first inequality follows from the convexity of $P(y)$. Therefore, when $\rho_{g_{1}}\rho_{g_{2}}\geq L_{P}^{2}$, the approximate secant inequality is satisfied. This establishes (3). ∎
2.3 On the Moreau Envelope Function Reformulation
This part explores key properties of the Moreau envelope function $v_{\gamma}(x,y)$ and the reformulation $({\rm VP})_{\gamma}$, as defined in (5) and analyzed in [29]. These properties form the foundation for the convergence analysis of the proposed algorithm.
Under Assumption 2, the solution mapping
$S_{\gamma}(x,y):=\mathrm{argmin}_{\theta\in Y}\left\{f(x,\theta)+g(x,\theta)+\frac{1}{2\gamma}\|\theta-y\|^{2}\right\}$ (8)
is well-defined and single-valued on $X\times\mathbb{R}^{m}$. We denote its unique solution by $\theta_{\gamma}^{*}(x,y)$. Below, we summarize sensitivity results for $v_{\gamma}(x,y)$ from Theorems 2, 3, and 5 of [29].
Proposition 1.
Suppose Assumptions 2 and 3 hold. For $\gamma\in(0,1/(2\rho_{f_{2}}+2\rho_{g_{2}}))$, the function $v_{\gamma}(x,y)$ is $(\rho_{v_{1}},\rho_{v_{2}})$-weakly convex on $X\times\mathbb{R}^{m}$ with $\rho_{v_{1}}\geq\rho_{f_{1}}+\rho_{g_{1}}$ and $\rho_{v_{2}}\geq 1/\gamma$. Additionally, $v_{\gamma}(x,y)$ is differentiable on $X\times\mathbb{R}^{m}$ with gradient
$\nabla v_{\gamma}(x,y)=\Big{(}\nabla_{x}f(x,\theta_{\gamma}^{*}(x,y))+\nabla_{x}g(x,\theta_{\gamma}^{*}(x,y)),\left({y-\theta_{\gamma}^{*}(x,y)}\right)/{\gamma}\Big{)}.$ (9)
In our algorithm’s convergence analysis, we require only the partial weak convexity of $v_{\gamma}(x,y)$, which differs from the full weak convexity used in [29]. This relaxed condition allows for a larger range of the regularization parameter $\gamma$ and smaller weak convexity constants of $v_{\gamma}(x,y)$, enabling larger step sizes in the algorithm. The partial weak convexity of $v_{\gamma}(x,y)$ is established below.
Proposition 2.
Suppose Assumptions 2 and 3 hold. For $\gamma\in(0,1/(\rho_{f_{2}}+\rho_{g_{2}}))$, the function $v_{\gamma}(x,y)$ is $\rho_{v_{1}}$-weakly convex with respect to $x$ on $X$ for any $y\in Y$ with $\rho_{v_{1}}\geq\rho_{f_{1}}+\rho_{g_{1}}$. It is also convex with respect to $y$ on $\mathbb{R}^{m}$ for any $x\in X$. Moreover, $v_{\gamma}(x,y)$ is differentiable with gradient given by (9) on $X\times\mathbb{R}^{m}$, and $\theta_{\gamma}^{*}(x,y)$ and $\nabla v_{\gamma}(x,y)$ are continuous on $X\times\mathbb{R}^{m}$.
Proof.
We begin by showing that $v_{\gamma}(x,y)$ is continuously differentiable on $X\times\mathbb{R}^{m}$. By Assumption 2, the function $f(x,\theta)+g(x,\theta)+\frac{1}{2\gamma}\|\theta-y\|^{2}$ is level-bounded in $\theta$ locally uniformly for any $(\bar{x},\bar{y})\in X\times\mathbb{R}^{m}$. Specifically, for any $c\in\mathbb{R}$, there exist a compact set $D$ and a neighborhood $Z$ of $(\bar{x},\bar{y})$ such that the level set $\{\theta\in\mathbb{R}^{m}\mid f(x,\theta)+g(x,\theta)+\frac{1}{2\gamma}\|\theta-y\|^{2}\leq c\}$ is contained in $D$ for all $(x,y)\in Z$. Using Corollary 4.3(ii) from [51], it follows that $v_{\gamma}(x,y)$ is locally Lipschitz continuous on $X\times\mathbb{R}^{m}$. Furthermore, Theorem 1.22, Theorem 4.1(ii) and Theorem 4.17 in [51] imply that $v_{\gamma}({x},{y})$ is differentiable on $X\times\mathbb{R}^{m}$ with its gradient given as in (9). The partial weak convexity property of $v_{\gamma}({x},{y})$ can be established using arguments similar to those in the proof of Theorem 2 in [29].
Next, we show that $\nabla v_{\gamma}({x},{y})$ is continuous on $X\times\mathbb{R}^{m}$ by showing the continuity of $\theta^{*}(x,y)$ on $X\times\mathbb{R}^{m}$. Let $\{(x^{k},y^{k})\}\subset X\times\mathbb{R}^{m}$ be any sequence such that $(x_{k},y_{k})\rightarrow(\bar{x},\bar{y})$. Since $f(x,\theta)+g(x,\theta)+\frac{1}{2\gamma}\|\theta-y\|^{2}$ is level-bounded in $\theta$ locally uniformly for $(\bar{x},\bar{y})\in X\times\mathbb{R}^{m}$, the sequence $\theta^{*}_{\gamma}(x_{k},y_{k})$ is bounded. Let $\bar{\theta}$ be an accumulation point of $\{\theta^{*}_{\gamma}(x_{k},y_{k})\}$. For any $\theta\in Y$, we have $f(x_{k},\theta^{*}_{\gamma}(x_{k},y_{k}))+g(x_{k},\theta^{*}_{\gamma}(x_{k},y_{k}))+\frac{1}{2\gamma}\|\theta^{*}_{\gamma}(x_{k},y_{k})-y_{k}\|^{2}\leq f(x_{k},\theta)+g(x_{k},\theta)+\frac{1}{2\gamma}\|\theta-y_{k}\|^{2}$. Taking the limit as $k\rightarrow\infty$, and using the continuity of $f$ and $g$, we obtain that for any $\theta\in Y$, $f(\bar{x},\bar{\theta})+g(\bar{x},\bar{\theta})+\frac{1}{2\gamma}\|\bar{\theta}-\bar{y}\|^{2}\leq f(\bar{x},\theta)+g(\bar{x},\theta)+\frac{1}{2\gamma}\|\theta-\bar{y}\|^{2}$. This implies $\bar{\theta}=\theta^{*}_{\gamma}(\bar{x},\bar{y})$. Therefore, $\theta^{*}_{\gamma}(x_{k},y_{k})\rightarrow\theta^{*}_{\gamma}(\bar{x},\bar{y})$, establishing the continuity of $\theta^{*}(x,y)$ on $X\times\mathbb{R}^{m}$. ∎
The solution mapping $\theta_{\gamma}^{*}(x,y)$ of the proximal problem defining $v_{\gamma}(x,y)$ exhibits Lipschitz continuity with respect to $y$, as established below.
Lemma 2.
Suppose Assumptions 2 and 3 hold. The mapping $\theta_{\gamma}^{*}(x,y)$ is $L_{\theta^{*}}$-Lipschitz continuous with respect to $y$ for any $x\in X$, with $L_{\theta^{*}}:=1+\sqrt{1-s/\gamma}$ where $s=\gamma/(\gamma L_{f_{y}}+1)^{2}$.
Proof.
For any $y_{1},y_{2}\in\mathbb{R}^{m}$, let $\theta^{*}_{1}:=\theta_{\gamma}^{*}(x,y_{1})$ and $\theta^{*}_{2}:=\theta_{\gamma}^{*}(x,y_{2})$ be optimal solutions to problem $\min_{\theta\in Y}\,\varphi(x,\theta)+\frac{1}{2\gamma}\|\theta-y\|^{2}$ with $y=y_{1}$ and $y_{2}$, respectively. From the first-order optimality conditions, we have
$20\in\nabla_{y}f(x,\theta^{*}_{i})+\partial_{y}g(x,\theta^{*}_{i})+(\theta^{*}_{i}-y_{i})/\gamma+\mathcal{N}_{Y}(\theta^{*}_{i}),\quad\mathrm{for}~{}i=1,$ 2 .
Since $\tilde{g}(x,y):=g(x,y)+\delta_{Y}(y)$ is convex with respect to $y$, we have
$\theta^{*}_{i}=\mathrm{Prox}_{s\tilde{g}(x,\cdot)}\left(\theta^{*}_{i}-s\left(\nabla_{y}f(x,\theta^{*}_{i})+(\theta^{*}_{i}-y_{i})/\gamma\right)\right),\quad\mathrm{for}~{}i=1,2,$ (10)
for any $s>0$. Using the nonexpansiveness of $\mathrm{Prox}_{s\tilde{g}(x,\cdot)}$, we have
$\|\theta^{*}_{1}-\theta^{*}_{2}\|$ (11)
$\leq$ $\big{\|}\left(\theta^{*}_{1}-s\left(\nabla_{y}f(x,\theta^{*}_{1})+(\theta^{*}_{1}-y_{1})/\gamma\right)\right)-\left(\theta^{*}_{2}-s\left(\nabla_{y}f(x,\theta^{*}_{2})+(\theta^{*}_{2}-y_{2})/\gamma\right)\right)\big{\|}.$
Since $f(x,\theta)+\frac{1}{2\gamma}\|\theta-y_{1}\|^{2}$ is $1/\gamma$-strongly convex with respect to $\theta$ on $Y$, we obtain
$\left\langle\nabla_{y}f(x,\theta^{*}_{1})+(\theta^{*}_{1}-y_{1})/\gamma-\nabla_{y}f(x,\theta^{*}_{2})-(\theta^{*}_{2}-y_{1})/\gamma,\theta^{*}_{1}-\theta^{*}_{2}\right\rangle\geq\frac{1}{\gamma}\|\theta^{*}_{1}-\theta^{*}_{2}\|^{2}.$
Let $s=\gamma/(\gamma L_{f_{y}}+1)^{2}$, then
$\big{\|}\left(\theta^{*}_{1}-s\left(\nabla_{y}f(x,\theta^{*}_{1})+(\theta^{*}_{1}-y_{1})/\gamma\right)\right)-\left(\theta^{*}_{2}-s\left(\nabla_{y}f(x,\theta^{*}_{2})+(\theta^{*}_{2}-y_{1})/\gamma\right)\right)\big{\|}^{2}$
$\leq$ $\left(1-s/\gamma\right)\|\theta^{*}_{1}-\theta^{*}_{2}\|^{2}.$
Combining this with (11), we derive
$\|\theta^{*}_{1}-\theta^{*}_{2}\|\leq\sqrt{1-s/\gamma}\big{\|}\theta^{*}_{1}-\theta^{*}_{2}\big{\|}+s/\gamma\cdot\|y_{1}-y_{2}\|,$ (12)
leading to the desired conclusion. ∎
Next, following the approaches in [29, 27], we provide a characterization of solution quality for problem $({\rm VP})_{\gamma}^{\epsilon}$. We begin by introducing a constraint qualification condition that will be used in the analysis.
Definition 1.
Let $(\bar{x},\bar{y})\in X\times Y$. We say that the extended no nonzero abnormal multiplier constraint qualification (ENNAMCQ) holds at $(\bar{x},\bar{y})$ for problem $({\rm VP})_{\gamma}^{\epsilon}$ if either $f(\bar{x},\bar{y})+g(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})<\epsilon$ or $f(\bar{x},\bar{y})+g(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})\geq\epsilon$ but
$0\notin\nabla f(\bar{x},\bar{y})+\partial g(\bar{x},\bar{y})-\nabla v_{\gamma}(\bar{x},\bar{y})+\mathcal{N}_{X\times Y}(\bar{x},\bar{y}).$ (13)
For any $\epsilon>0$, as shown in Proposition 17 of [29], ENNAMCQ holds for problem $({\rm VP})_{\gamma}^{\epsilon}$ if the lower-level problem is strictly convex with respect to $y$.
Proposition 3.
Suppose that Assumptions 2 and 3 hold and $\gamma>0$. Then for any $\epsilon>0$ and $(\bar{x},\bar{y})\in X\times Y$, problem $({\rm VP})_{\gamma}^{\epsilon}$ satisfies ENNAMCQ at $(\bar{x},\bar{y})$, provided that the lower level problem is strictly convex for $\bar{x}$, i.e., $\varphi(\bar{x},y)+\delta_{Y}(y)$ is strictly convex with respect to $y$.
Having established the constraint qualification condition, we now define the concept of a stationary point for problem $({\rm VP})_{\gamma}^{\epsilon}$. This serves as a candidate for optimal solutions of the problem under the ENNAMCQ.
Definition 2.
Let $(\bar{x},\bar{y})$ be a feasible solution of problem $({\rm VP})_{\gamma}^{\epsilon}$ with $\epsilon>0$. We say that $(\bar{x},\bar{y})$ is a stationary/KKT point of problem $({\rm VP})_{\gamma}^{\epsilon}$ if there exists a multiplier $\lambda\geq 0$ such that
$\begin{cases}0\in\nabla F(\bar{x},\bar{y})+\lambda\left(\nabla f(\bar{x},\bar{y})+\partial g(\bar{x},\bar{y})-\nabla v_{\gamma}(\bar{x},\bar{y})\right)+\mathcal{N}_{X\times Y}(\bar{x},\bar{y}),\\ f(\bar{x},\bar{y})+g(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})\leq\epsilon,\qquad\lambda\left(f(\bar{x},\bar{y})+g(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})-\epsilon\right)=0.\end{cases}$
Using 3, following [29, 27], we have the following necessary optimality condition.
Theorem 1.
Assume Assumptions 2 and 3 hold and $\gamma>0$. Let $(\bar{x},\bar{y})$ be a local optimal solution to problem ${\rm(VP)}_{\gamma}^{\epsilon}$ with $\epsilon>0$, and the lower level problem is strictly convex with respect to $y$ for $\bar{x}$. Then $(\bar{x},\bar{y})$ is a KKT point of problem ${\rm(VP)}_{\gamma}^{\epsilon}$.
3 Algorithm design
In this section, we propose a new gradient-type algorithm, Alternating Gradient-type algorithm with Inexact Lower-level Solutions (AGILS), to solve ${\rm(VP)}_{\gamma}^{\epsilon}$. Given that $f(x,y)+g(x,y)-v_{\gamma}(x,y)$ is always nonnegative, a natural approach is to consider the following penalized problem,
$\min_{x\in X,y\in Y}\psi_{p_{k}}(x,y):=\frac{1}{p_{k}}F(x,y)+f(x,y)+g(x,y)-v_{\gamma}(x,y),$
where $p_{k}$ is the penalty parameter at iteration $k$. Due the nonsmoothness of $g(x,y)$ with respect to $y$ and the availability of its proximal operator, it is natural to adopt a proximal alternating linearization method [52] to construct the algorithm. Given the current iterate $(x^{k},y^{k})$, linearizing the smooth components of $\psi_{p_{k}}(x,y)$ requires evaluating gradients, including $v_{\gamma}(x,y)$ as provided in 1. Calculating the exact gradient $\nabla v_{\gamma}(x^{k},y^{k})$ requires solving the proximal lower-level problem in (8) to obtain $\theta_{\gamma}^{*}(x^{k},y^{k})$. However, this solution typically does not have a closed-form expression, and solving it requires an iterative optimization solver, which can be computationally expensive. To address this, we propose an inexact gradient approximation by substituting $\theta_{\gamma}^{*}$ with an inexact estimate $\theta^{k}$.
The proposed algorithm alternates between updating $y$ and $x$ while refining the inexact approximation of $\theta_{\gamma}^{*}(x^{k},y^{k})$. Given a current iterate $(x^{k},y^{k})$ and $\theta^{k}$ for each $k=0,1,\ldots$, the next iterate $(x^{k+1},y^{k+1})$ and $\theta^{k+1}$ are updated as follows.
Update $y$ (Fixing $x$). We compute the update direction for $y$,
$d_{y}^{k}:=\frac{1}{p_{k}}\nabla_{y}F\left(x^{k},y^{k}\right)+\nabla_{y}f\left(x^{k},y^{k}\right)-\frac{y^{k}-\theta^{k}}{\gamma},$ (14)
where $(y^{k}-\theta^{k})/{\gamma}$ is an inexact approximation of $\nabla_{y}v_{\gamma}(x^{k},y^{k})=({y^{k}}-\theta_{\gamma}^{*}(x^{k},y^{k}))/\gamma$ (see (9)). Using $d_{y}^{k}$, the variable $y$ is updated as
$y^{k+1}=\underset{y\in Y}{\mathrm{argmin}}~{}\left\langle d_{y}^{k},y-y^{k}\right\rangle+g(x^{k},y)+\frac{1}{2\beta_{k}}\left\|y-y^{k}\right\|^{2},$ (15)
where $\beta_{k}>0$ is the stepsize. This can also be expressed using the proximal operator
$y^{k+1}=\operatorname{Prox}_{\beta_{k}\tilde{g}(x^{k},\cdot)}\left(y^{k}-\beta_{k}d^{k}_{y}\right),$ (16)
where $\tilde{g}(x,y):=g(x,y)+\delta_{Y}(y)$.
Find inexact approximation $\theta^{k+1/2}$. After updating $y^{k+1}$, we compute an inexact approximation $\theta^{k+1/2}$ for $\theta^{*}(x^{k},y^{k+1})$ by solving
$\theta^{k+1/2}\approx\underset{y\in Y}{\mathrm{argmin}}\left\{f(x^{k},\theta)+g(x^{k},\theta)+\frac{1}{2\gamma}\left\|\theta-y^{k+1}\right\|^{2}\right\},$ (17)
subject to the inexact criterion
$\mathcal{G}(\theta^{k+1/2},x^{k},y^{k+1})\leq s_{k},$ (18)
where $s_{k}$ controls the inexactness, and satisfies $\sum_{k=0}^{\infty}s_{k}^{2}<\infty$. The prox-gradient residual $\mathcal{G}(\theta,x,y)$ quantifies the inexactness and is defined as
$\mathcal{G}(\theta,x,y):=\left\|\theta-\operatorname{Prox}_{\eta\tilde{g}\left(x,\cdot\right)}\left(\theta-\eta\left(\nabla_{y}f\left(x,\theta\right)+(\theta-y)/{\gamma}\right)\right)\right\|,$
where $\eta>0$. Note that $\mathcal{G}(\theta,x,y)=0$ if and only if $\theta=\theta^{*}(x,y)$.
Update $x$ (Fixing $y$). We compute the update direction for $x$,
$d_{x}^{k}:=\frac{1}{p_{k}}\nabla_{x}F\left(x^{k},y^{k+1}\right)+\nabla_{x}\varphi\left(x^{k},y^{k+1}\right)-\nabla_{x}f\left(x^{k},\theta^{k+1/2}\right)-\nabla_{x}g\left(x^{k},\theta^{k+1/2}\right),$ (19)
where $\nabla_{x}f\left(x^{k},\theta^{k+1/2}\right)-\nabla_{x}g\left(x^{k},\theta^{k+1/2}\right)$ is an approximation to $\nabla_{x}v_{\gamma}(x^{k},y^{k})=\nabla_{x}f(x^{k},\theta^{*}(x^{k},y^{k}))-\nabla_{x}g(x^{k},\theta^{*}(x^{k},y^{k}))$ (see (9)). Using the direction $d_{x}^{k}$, we update $x$ with stepsize $\alpha_{k}>0$ as
$x^{k+1}=\operatorname{Proj}_{X}\left(x^{k}-\alpha_{k}d_{x}^{k}\right).$ (20)
Find inexact approximation $\theta^{k+1}$. After updating $x^{k+1}$, we compute a new inexact approximation $\theta^{k+1}$ for $\theta^{*}(x^{k+1},y^{k+1})$ satisfying the inexact criterion
$\mathcal{G}(\theta^{k+1},x^{k+1},y^{k+1})\leq s_{k+1}.$ (21)
Update the penalty parameter $p_{k+1}$. To ensure the feasibility of the constraint $\varphi(x,y)-v_{\gamma}(x,y)\leq\epsilon$, we update the penalty parameter $p_{k}$ at each iteration. Evaluating the constraint violation requires computing $v_{\gamma}(x,y)$, which involves solving the proximal lower-level problem in (8). Instead, we use $\theta^{k+1}$ to update the penalty parameter. Specifically, we compute an approximation to the constraint violation as
$t^{k+1}=\max\left\{\varphi(x^{k+1},y^{k+1})-\varphi(x^{k+1},\theta^{k+1})-\frac{1}{2\gamma}\|\theta^{k+1}-y^{k+1}\|^{2}-\epsilon,0)\right\}.$ (22)
The penalty parameter is then updated as
$p_{k+1}=\begin{cases}p_{k}+\varrho_{p},&\text{if}\quad\left\|(x^{k+1},y^{k+1})-(x^{k},y^{k})\right\|<c_{p}\min\left\{1/p_{k},{t_{k+1}}\right\},\\ p_{k},&\text{otherwise},\end{cases}$ (23)
where $\varrho_{p}$ and $c_{p}$ are positive constants controlling the penalty update.
With these steps, we are now ready to present the proposed AGILS in 1.
Algorithm 1 Alternating Gradient-type algorithm with Inexact Lower-level Solutions (AGILS)
1:  Input: Initial iterates $x^{0},y^{0},\theta^{0}$, stepsizes $\alpha_{k},\beta_{k}$, relaxation parameter $\epsilon$, proximal parameter $\gamma$, penalty parameters $p_{0}$, $\varrho_{p}$, $c_{p}$, inexact parameters $\eta$, $s_{k}$, tolerance $tol$.
[…]
3:     Construct the direction $d_{y}^{k}$ according to (14), and update $y^{k+1}$ as
$y^{k+1}=\operatorname{Prox}_{\beta_{k}\tilde{g}(x^{k,\cdot})}\left(y^{k}-\beta_{k}d^{k}_{y}\right).$
4:     Find an inexact solution $\theta^{k+1/2}$ that satisfies the inexactness criterion (18).
5:     Construct the direction $d_{x}^{k}$ according to (19), and update $x^{k+1}$ as
$x^{k+1}=\operatorname{Proj}_{X}\left(x^{k}-\alpha_{k}d_{x}^{k}\right).$
6:     Find an inexact solution $\theta^{k+1}$ that satisfies the inexactness criterion (21).
7:     Stopping test. Compute $t^{k+1}$ according to (22). Stop when $k\geq 1$ and $\max\{s_{k},\|(x^{k+1},y^{k+1})-(x^{k},y^{k})\|,t^{k+1}\}\leq tol$.
8:     Penalty parameter update. Set $p_{k+1}$ according to (23).
9:  end for
10:  return  $(x^{k+1},y^{k+1})$
4 Convergence properties
In this section, we analyze the convergence of the proposed AGILS for solving $({\rm VP})_{\gamma}^{\epsilon}$. Specifically, we establish the subsequential convergence of AGILS to the KKT points of the $({\rm VP})_{\gamma}^{\bar{\epsilon}}$ for some $\bar{\epsilon}\leq\epsilon$.
4.1 Preliminary results for the analysis
We begin by considering the following lemma, which provides an estimate for the distance between the approximation $\theta$ and the exact solution $\theta^{*}(x,y)$ of the proximal lower-level problem, based on the prox-gradient residual $\mathcal{G}(\theta,x,y)$ in the inexact criteria (18) and (21). This estimate is derived from the $1/\gamma$-strongly convexity and $1/\gamma+L_{f}$-smoothness of the component $f(x,\theta)+\|\theta-y\|^{2}/(2\gamma)$ with respect to $\theta$ in the objective of the proximal lower level problem, see, e.g., [53, Theorem 3.4 and 3.5] or [54, Proposition 2.2].
Lemma 3.
Suppose Assumptions 2 and 3 hold, and let $\gamma>0$, $\eta>0$. Then, there exists $C_{\eta}>0$ such that for any $(x,y)\in X\times{\mathbb{R}}^{m}$, and $\theta\in{\mathbb{R}}^{m}$, it holds that
$\|\theta-\theta^{*}(x,y)\|\leq C_{\eta}\left\|\theta-\operatorname{Prox}_{\eta\tilde{g}\left(x,\cdot\right)}\left(\theta-\eta\left(\nabla_{y}f\left(x,\theta\right)+(\theta-y)/{\gamma}\right)\right)\right\|,$
where $\theta^{*}(x,y):=\arg\min_{\theta\in Y}\{\varphi(x,\theta)+\frac{1}{2\gamma}\|\theta-y\|^{2}\}$.
The next lemma provides important inequalities involving $v_{\gamma}$ and plays a crucial role in proving the decreasing property of the proposed algorithm.
Lemma 4.
Suppose Assumptions 2 and 3 hold. Let $\gamma\in(0,1/(\rho_{f_{2}}+\rho_{g_{2}}))$ and $(\bar{x},\bar{y})\in X\times{\mathbb{R}}^{m}$. Then, for any $(x,y)\in X\times{\mathbb{R}}^{m}$, the following inequalities holds,
$-v_{\gamma}(x,\bar{y})$ $\leq-v_{\gamma}(\bar{x},\bar{y})-\langle\nabla_{x}v_{\gamma}(\bar{x},\bar{y}),x-\bar{x}\rangle+\frac{\rho_{f_{1}}+\rho_{g_{1}}}{2}\|\bar{x}-x\|^{2},$ (24)
$-v_{\gamma}(\bar{x},y)$ $\leq-v_{\gamma}(\bar{x},\bar{y})-\langle\nabla_{y}v_{\gamma}(\bar{x},\bar{y}),y-\bar{y}\rangle.$ (25)
Proof.
According to Proposition 2, for any $\bar{y}\in{\mathbb{R}}^{m}$, $v_{\gamma}(x,\bar{y})$ is a $\rho_{f_{1}}+\rho_{g_{1}}$-weakly convex function with respect to $x$ on $X$, that is, $v_{\gamma}(x,\bar{y})+\frac{\rho_{f_{1}}+\rho_{g_{1}}}{2}\|x\|^{2}$ is convex on $X$. Then we can obtain (24) immediately. The second inequality (25) follows from Proposition 2 that for any $\bar{x}\in X$, $v_{\gamma}(\bar{x},y)$ is convex with respect to $y$. ∎
We now proceed to establish a decreasing property for the proposed method based on the following merit function:
$\tilde{\psi}_{p_{k}}(x,y):=\frac{1}{p_{k}}(F(x,y)-\underline{F})+f(x,y)+g(x,y)-v_{\gamma}(x,y).$
Lemma 5.
Suppose Assumptions 1, 2 and 3 hold. Let $\gamma\in(0,1/(\rho_{f_{2}}+\rho_{g_{2}}))$, $\alpha_{k}\in[0,2/(L_{\psi_{x,k}}+c_{\alpha})]$ and $\beta_{k}\in[0,2/(L_{\psi_{y,k}}+c_{\beta})]$, where $L_{\psi_{x,k}}=L_{F_{x}}/p_{k}+L_{f_{x}}+L_{g_{1}}+\rho_{f_{1}}+\rho_{g_{1}}$, $L_{\psi_{y,k}}=L_{F_{y}}/p_{k}+L_{f_{y}}$ and $c_{\alpha},c_{\beta}$ are positive constants. Then, there exists $C_{\eta}>0$ such that the sequence $(x^{k},y^{k})$ generated by AGILS (1) satisfies the following inequality,
$\tilde{\psi}_{p_{k+1}}\left(x^{k+1},y^{k+1}\right)\leq$ $\tilde{\psi}_{p_{k}}\left(x^{k},y^{k}\right)+\left(\frac{1}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}+\frac{1}{c_{\beta}\gamma^{2}}\right)C_{\eta}^{2}s_{k}^{2}$ (26)
$-\frac{c_{\alpha}}{4}\left\|x^{k+1}-x^{k}\right\|^{2}-\frac{c_{\beta}}{4}\left\|y^{k+1}-y^{k}\right\|^{2}.$
Proof.
First, by the update rule for $y^{k+1}$ given in (15), we have
$g(x^{k},y^{k+1})+\langle d_{y}^{k},y^{k+1}-y^{k}\rangle+\frac{1}{\beta_{k}}\|y^{k+1}-y^{k}\|^{2}\leq g(x^{k},y^{k}).$ (27)
According to the assumptions, with $x^{k}\in X$, we have that $\nabla_{y}F(x^{k},y)$ and $\nabla_{y}f(x^{k},y)$ are $L_{F_{y}}$- and $L_{f_{y}}$-Lipschitz continuous with respect to variable $y$ on $Y$, respectively, and combine this with (25), we have
$(\tilde{\psi}_{p_{k}}-g)(x^{k},y^{k+1})$ (28)
$\leq$ $(\tilde{\psi}_{p_{k}}-g)(x^{k},y^{k})+\langle\nabla_{y}(\tilde{\psi}_{p_{k}}-g)(x^{k},y^{k}),y^{k+1}-y^{k}\rangle+\frac{L_{\psi_{y,k}}}{2}\|y^{k+1}-y^{k}\|^{2},$
where $L_{\psi_{y,k}}=L_{F_{y}}/p_{k}+L_{f_{y}}$. Combining (27) and (28) yields
$\tilde{\psi}_{p_{k}}\left(x^{k},y^{k+1}\right)\leq$ $\tilde{\psi}_{p_{k}}\left(x^{k},y^{k}\right)+\langle\nabla_{y}(\tilde{\psi}_{p_{k}}-g)(x^{k},y^{k})-d_{y}^{k},y^{k+1}-y^{k}\rangle$ (29)
$-\left(\frac{1}{\beta_{k}}-\frac{L_{\psi_{y,k}}}{2}\right)\left\|y^{k+1}-y^{k}\right\|^{2}.$
Using the formula of the gradient of $v_{\gamma}$ given in (9) and the construction of $d_{y}^{k}$ in (14), and 3 and the inexact condition (18), we have
$\left\|\nabla_{y}(\tilde{\psi}_{p_{k}}-g)(x^{k},y^{k})-d_{y}^{k}\right\|$ $=\|\nabla_{y}v_{\gamma}(x^{k},y^{k})+(y^{k}-\theta^{k})/\gamma\|$ (30)
$=\frac{1}{\gamma}\left\|\theta^{*}(x^{k},y^{k})-\theta^{k}\right\|\leq\frac{C_{\eta}}{\gamma}s_{k}.$
Combining the above inequality with (29) gives us
$\tilde{\psi}_{p_{k}}\left(x^{k},y^{k+1}\right)\leq\,\tilde{\psi}_{p_{k}}\left(x^{k},y^{k}\right)-\left(\frac{1}{\beta_{k}}-\frac{L_{\psi_{y,k}}}{2}-\frac{c_{\beta}}{4}\right)\left\|y^{k+1}-y^{k}\right\|^{2}+\frac{C_{\eta}^{2}}{c_{\beta}\gamma^{2}}s_{k}^{2}.$ (31)
Next, considering the update rule for $x^{k+1}$ given in (20), we have
$\langle d_{x}^{k},x^{k+1}-x^{k}\rangle+\frac{1}{\alpha_{k}}\|x^{k+1}-x^{k}\|^{2}\leq 0.$ (32)
By the assumptions, with $y^{k+1}\in Y$, we have that $\nabla_{x}F(x,y^{k+1})$, $\nabla_{x}f(x,y^{k+1})$ and $\nabla_{x}g(x,y^{k+1})$ are $L_{F_{x}}$-, $L_{f_{x}}$-, and $L_{g_{1}}$-Lipschitz continuous with respect to variable $x$ on $X$, respectively, and combine this with (24), we have
$\tilde{\psi}_{p_{k}}(x^{k+1},y^{k+1})\leq\tilde{\psi}_{p_{k}}(x^{k},y^{k+1})+\langle\nabla_{x}\tilde{\psi}_{p_{k}}(x^{k},y^{k+1}),x^{k+1}-x^{k}\rangle+\frac{L_{\psi_{x,k}}}{2}\|x^{k+1}-x^{k}\|^{2},$ (33)
where $L_{\psi_{x,k}}=L_{F_{x}}/p_{k}+L_{f_{x}}+L_{g_{1}}+\rho_{f_{1}}+\rho_{g_{1}}$. Combining (32) and (33) yields
$\tilde{\psi}_{p_{k}}\left(x^{k+1},y^{k+1}\right)\leq$ $\tilde{\psi}_{p_{k}}\left(x^{k},y^{k+1}\right)+\langle\nabla_{x}\tilde{\psi}_{p_{k}}(x^{k},y^{k+1})-d_{x}^{k},x^{k+1}-x^{k}\rangle$ (34)
$-\left(\frac{1}{\alpha_{k}}-\frac{L_{\psi_{x,k}}}{2}\right)\left\|x^{k+1}-x^{k}\right\|^{2}.$
Using the formula of the gradient of $v_{\gamma}$ given in (9) and the construction of $d_{x}^{k}$ in (19), and the $L_{f_{x}}$- and $L_{g_{2}}$-Lipschitz continuity of $\nabla_{x}f(x^{k},y)$ and $\nabla_{x}g(x^{k},y)$ with respect to variable $y$ on $Y$, we have
$\left\|\nabla_{x}\tilde{\psi}_{p_{k}}(x^{k},y^{k+1})-d_{x}^{k}\right\|$ $=\left\|\nabla_{x}\varphi(x^{k},\theta^{*}(x^{k},y^{k+1}))-\nabla_{x}\varphi(x^{k},\theta^{k+1/2})\right\|$ (35)
$\leq(L_{f_{x}}+L_{g_{2}})\left\|\theta^{*}(x^{k},y^{k+1})-\theta^{k+1/2}\right\|\leq(L_{f_{x}}+L_{g_{2}})C_{\eta}s_{k},$
where the last inequality follows from 3 and the inexact criterion (21). Combining the above inequality with (34) gives us
$\tilde{\psi}_{p_{k}}\left(x^{k+1},y^{k+1}\right)\leq$ $\tilde{\psi}_{p_{k}}\left(x^{k},y^{k+1}\right)-\left(\frac{1}{\alpha_{k}}-\frac{L_{\psi_{x,k}}}{2}-\frac{c_{\alpha}}{4}\right)\left\|x^{k+1}-x^{k}\right\|^{2}$ (36)
$+\frac{1}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}C_{\eta}^{2}s_{k}^{2}.$
Combining (31) and (36) yields the following decreasing property,
$\tilde{\psi}_{p_{k}}\left(x^{k+1},y^{k+1}\right)\leq$ $\tilde{\psi}_{p_{k}}\left(x^{k},y^{k}\right)+\frac{1}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}C_{\eta}^{2}s_{k}^{2}+\frac{C_{\eta}^{2}}{c_{\beta}\gamma^{2}}s_{k}^{2}$ (37)
$-\left(\frac{1}{\alpha_{k}}-\frac{L_{\psi_{x,k}}}{2}-\frac{c_{\alpha}}{4}\right)\left\|x^{k+1}-x^{k}\right\|^{2}-\left(\frac{1}{\beta_{k}}-\frac{L_{\psi_{y,k}}}{2}-\frac{c_{\beta}}{4}\right)\left\|y^{k+1}-y^{k}\right\|^{2}.$
Because $\alpha_{k}\in[0,2/(L_{\psi_{x,k}}+c_{\alpha})]$ and $\beta_{k}\in[0,2/(L_{\psi_{y,k}}+c_{\beta})]$, and $\tilde{\psi}_{p_{k+1}}(x^{k+1},y^{k+1})\leq\tilde{\psi}_{p_{k}}(x^{k+1},y^{k+1})$ following from the non-decreasing property of the penalty parameter $p_{k}$ and the non-negative property of $F-\underline{F}$, we can obtain the conclusion from (37). ∎
4.2 Convergence to stationary points
In this part, we establish the sequential convergence properties of the AGILS, described in 1. Throughout this subsection, we assume that Assumptions 1, 2, and 3 hold, and $\gamma\in(0,1/(\rho_{f_{2}}+\rho_{g_{2}}))$. The sequences $\{(x^{k},y^{k})\}$ and $\{p_{k}\}$ are generated by the algorithm, where the step sizes satisfy $\alpha_{k}\in[\underline{\alpha},2/(L_{\psi_{x,k}}+c_{\alpha})]$ and $\beta_{k}\in[\underline{\beta},2/(L_{\psi_{y,k}}+c_{\beta})]$, where $L_{\psi_{x,k}}=L_{F_{x}}/p_{k}+L_{f_{x}}+L_{g_{1}}+\rho_{f_{1}}+\rho_{g_{1}}$, $L_{\psi_{y,k}}=L_{F_{y}}/p_{k}+L_{f_{y}}$ and $\underline{\alpha},\underline{\beta},c_{\alpha},c_{\beta}$ are positive constants, and $s_{k}$ satisfies $\sum_{k=0}^{\infty}s_{k}^{2}<\infty$. Our analysis demonstrates the sequential convergence of $\{(x^{k},y^{k})\}$ to a stationary point of $({\rm VP})_{\gamma}^{\bar{\epsilon}}$ for some $\bar{\epsilon}\leq\epsilon$.
We first show that the differences between consecutive iterates are square summable.
Lemma 6.
The sequence $\left\{(x^{k},y^{k})\right\}$ satisfies
$\sum\limits_{k=1}^{{\infty}}\left\|(x^{k+1},y^{k+1})-(x^{k},y^{k})\right\|^{2}<\infty,\quad\text{and}\quad\lim\limits_{k\to\infty}\left\|(x^{k+1},y^{k+1})-(x^{k},y^{k})\right\|=0.$
Proof.
Summing 26 from Lemma 5 over $k=0,1,\dots,K-1$ yields,
$\tilde{\psi}_{p_{K}}\left(x^{K},y^{K}\right)-\tilde{\psi}_{p_{0}}\left(x^{0},y^{0}\right)\leq$ $-\sum\limits_{k=0}^{K-1}\left(\frac{c_{\alpha}}{4}\left\|x^{k+1}-x^{k}\right\|^{2}+\frac{c_{\beta}}{4}\left\|y^{k+1}-y^{k}\right\|^{2}\right)$ (38)
$+\sum\limits_{k=0}^{K-1}\left(\frac{1}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}+\frac{1}{c_{\beta}\gamma^{2}}\right)C_{\eta}^{2}s_{k}^{2}.$
Since $\{s_{k}\}$ is square summable, it follows that
$\sum\limits_{k=0}^{\infty}\left(\frac{1}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}+\frac{1}{c_{\beta}\gamma^{2}}\right)C_{\eta}^{2}s_{k}^{2}<\infty.$
Additionally, because $\tilde{\psi}_{p_{K}}\left(x^{K},y^{K}\right)\geq 0$, by taking $K\rightarrow\infty$ in (38), we have
$\quad\,\sum\limits_{k=0}^{\infty}\left(\frac{c_{\alpha}}{4}\left\|x^{k+1}-x^{k}\right\|^{2}+\frac{c_{\beta}}{4}\left\|y^{k+1}-y^{k}\right\|^{2}\right)$ (39)
$\leq\tilde{\psi}_{p_{0}}\left(x^{0},y^{0}\right)+\sum\limits_{k=0}^{\infty}\left(\frac{1}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}+\frac{1}{c_{\beta}\gamma^{2}}\right)C_{\eta}^{2}s_{k}^{2}<\infty,$
which proves the result. ∎
We now present the main result, showing that every accumulation point of sequence ${(x^{k},y^{k})}$ is a KKT point of the problem $({\rm VP})_{\gamma}^{\bar{\epsilon}}$ for some $\bar{\epsilon}\leq\epsilon$.
Theorem 2.
Let $\epsilon>0$, if the sequence $\{p_{k}\}$ is bounded, then every accumulation point $(\bar{x},\bar{y})$ of $\{(x^{k},y^{k})\}$ is a KKT point of $({\rm VP})_{\gamma}^{\bar{\epsilon}}$ for some $\bar{\epsilon}\leq\epsilon$.
Proof.
The proof builds upon techniques used in the proof of Theorem 1 in [27]. Let $(\bar{x},\bar{y})$ be an accumulation point of the sequence $\{(x^{k},y^{k})\}$, and let $\{(x^{k_{j}},y^{k_{j}})\}$ be the corresponding subsequence such that $(x^{k_{j}},y^{k_{j}})\rightarrow(\bar{x},\bar{y})$ as $k_{j}\rightarrow\infty$. Since $\{p_{k}\}$ is bounded, there exists $k_{0}>0$ such that $p_{k}=\bar{p}:=p_{k_{0}}$ for any $k\geq k_{0}$. From 6, we know that
$\lim\limits_{k\to\infty}\left\|(x^{k+1},y^{k+1})-(x^{k},y^{k})\right\|=0,$ (40)
which implies that for sufficiently large $k$, the inequality $\left\|(x^{k+1},y^{k+1})-(x^{k},y^{k})\right\|\leq c_{p}/p_{k}$ holds. According to the update strategy of $p_{k}$ given in (23), we have, for large $k$, that $t^{k}<c_{p}^{-1}\|(x^{k},y^{k})-(x^{k-1},y^{k-1})\|$ and thus $t^{k}\rightarrow 0$. Consequently, from the definition of $t^{k}$, we obtain
$\max\left\{\varphi(x^{k},y^{k})-\varphi(x^{k},\theta^{k})-\frac{1}{2\gamma}\|\theta^{k}-y^{k}\|^{2}-\epsilon,0)\right\}\rightarrow 0.$ (41)
Next, using the update rules for $x^{k+1}$ and $y^{k+1}$ given in (20) and (15), and the expressions of $d_{x}^{k}$ and $d^{k}_{y}$, we have
$\xi_{x}^{k}$ $\in\nabla_{x}F(x^{k+1},y^{k+1})+p_{k}\left(\nabla_{x}\varphi(x^{k+1},y^{k+1})-\nabla_{x}v_{\gamma}(x^{k+1},y^{k+1})\right)+\mathcal{N}_{X}(x^{k+1}),$ (42)
$\xi_{y}^{k}$ $\in\nabla_{y}F(x^{k},y^{k+1})+p_{k}\left(\partial_{y}\varphi(x^{k},y^{k+1})-\nabla v_{\gamma}(x^{k},y^{k+1})\right)+\mathcal{N}_{Y}(y^{k+1}),$
where $\xi_{x}^{k}\in{\mathbb{R}}^{n}$ and $\xi_{y}^{k}\in{\mathbb{R}}^{m}$ are given by
$\xi_{x}^{k}$ $:=p_{k}\left(\nabla_{x}\tilde{\psi}_{p_{k}}\left(x^{k+1},y^{k+1}\right)-d_{x}^{k}\right)-\frac{p_{k}}{\alpha_{k}}\left(x^{k+1}-x^{k}\right),$ (43)
$\xi_{y}^{k}$ $:=p_{k}\left(\nabla_{y}\left(\tilde{\psi}_{p_{k}}-g\right)\left(x^{k},y^{k+1}\right)-d_{y}^{k}\right)-\frac{p_{k}}{\beta_{k}}\left(y^{k+1}-y^{k}\right).$
We now show that $\xi_{x}^{k_{j}},\xi_{y}^{k_{j}}\rightarrow 0$ as $k_{j}\rightarrow\infty$. First, we have
$\|\nabla_{x}\tilde{\psi}_{p_{k}}\left(x^{k+1},y^{k+1}\right)-d_{x}^{k}\|$ (44)
$\leq$ $\|\nabla_{x}\tilde{\psi}_{p_{k}}\left(x^{k+1},y^{k+1}\right)-\nabla_{x}\tilde{\psi}_{p_{k}}\left(x^{k},y^{k+1}\right)\|+\|\nabla_{x}\tilde{\psi}_{p_{k}}\left(x^{k},y^{k+1}\right)-d_{x}^{k}\|$
$\leq$ $\|\nabla_{x}\tilde{\psi}_{p_{k}}\left(x^{k+1},y^{k+1}\right)-\nabla_{x}\tilde{\psi}_{p_{k}}\left(x^{k},y^{k+1}\right)\|+(L_{f_{x}}+L_{g_{2}})C_{\eta}s_{k},$
where the last inequality follows from (35). Since $p_{k}=\bar{p}$ for all $k\geq k_{0}$, and assuming that $\nabla_{x}F,\nabla_{x}f$ and $\nabla_{x}g$ are continuous on $X\times Y$, and $\nabla_{x}v_{\gamma}$ is continuous on $X\times Y$ (Proposition 2), we conclude that $\nabla_{x}\tilde{\psi}_{p_{k}}=\nabla_{x}\tilde{\psi}_{\bar{p}}$ is continuous on $X\times Y$. As $\lim_{k_{j}\to\infty}(x^{k_{j}+1},y^{k_{j}+1})=\lim_{k_{j}\to\infty}(x^{k_{j}},y^{k_{j}})=(\bar{x},\bar{y})$ and $s_{k}\rightarrow 0$, it follows that $\nabla_{x}\tilde{\psi}_{p_{k_{j}}}(x^{k_{j}+1},y^{k_{j}+1})-d_{x}^{k_{j}}\rightarrow 0$ as $k_{j}\rightarrow\infty$. Combining this with (40) and (43) and the fact that $\alpha_{k}\geq\underline{\alpha}>0$, we obtain that $\xi_{x}^{k_{j}}\rightarrow 0$ as $k_{j}\rightarrow\infty$. A similar argument can be applied to $\xi_{y}^{k}$. From (30), we have
$\|\nabla_{y}\left(\tilde{\psi}_{p_{k}}-g\right)\left(x^{k},y^{k+1}\right)-d_{y}^{k}\|$ (45)
$\leq$ $\|\nabla_{y}\left(\tilde{\psi}_{p_{k}}-g\right)\left(x^{k},y^{k+1}\right)-\nabla_{y}\left(\tilde{\psi}_{p_{k}}-g\right)\left(x^{k},y^{k}\right)\|+{C_{\eta}}s_{k}/\gamma.$
Since $p_{k}=\bar{p}$ for all $k\geq k_{0}$, and under the assumption that $\nabla_{y}F$ and $\nabla_{y}f$ are continuous on $X\times Y$, and $\nabla_{y}v_{\gamma}$ is continuous on $X\times Y$ (Proposition 2), we have that $\nabla_{y}(\tilde{\psi}_{p_{k}}-g)=\nabla_{y}(\tilde{\psi}_{\bar{p}}-g)$ is continuous on $X\times Y$. As $\lim_{k_{j}\to\infty}(x^{k_{j}},y^{k_{j}+1})=\lim_{k_{j}\to\infty}(x^{k_{j}},y^{k_{j}})=(\bar{x},\bar{y})$ and $s_{k}\rightarrow 0$, it follows that $\nabla_{y}(\tilde{\psi}_{p_{k_{j}}}-g)(x^{k_{j}},y^{k_{j}+1})-d_{y}^{k_{j}}\rightarrow 0$ as $k_{j}\rightarrow\infty$. Combining this with (40), (45) and the fact that $\beta_{k}\geq\underline{\beta}>0$, we conclude that $\xi_{y}^{k_{j}}\rightarrow 0$ as $k_{j}\rightarrow\infty$.
Now, taking the limit as $k=k_{j}\rightarrow\infty$ in (42), since $\xi_{x}^{k_{j}},\xi_{y}^{k_{j}}\rightarrow 0$, $p_{k}=\bar{p}$ for all $k\geq k_{0}$, $\lim_{k_{j}\to\infty}(x^{k_{j}+1},y^{k_{j}+1})=\lim_{k_{j}\to\infty}(x^{k_{j}},y^{k_{j}+1})=\lim_{k_{j}\to\infty}(x^{k_{j}},y^{k_{j}})=(\bar{x},\bar{y})$, $\nabla F$, $\nabla v_{\gamma}$, and $\nabla_{x}\varphi$ are continuous on $X\times Y$, $\partial_{y}\varphi$ and $\mathcal{N}_{X\times Y}$ are outer semicontinuous on $X\times Y$, and $\partial\varphi=\{\nabla_{x}\varphi\}\times\partial_{y}\varphi$ on $X\times Y$, we obtain
$0\in\nabla F(\bar{x},\bar{y})+\bar{p}\left(\partial\varphi(\bar{x},\bar{y})-\nabla v_{\gamma}(\bar{x},\bar{y})\right)+\mathcal{N}_{X\times Y}(\bar{x},\bar{y}).$ (46)
Finally, since $s_{k}\rightarrow 0$, we have $\|\theta^{k}-\theta^{*}(x^{k},y^{k})\|\rightarrow 0$, and thus $\varphi(x^{k_{j}},\theta^{k_{j}})-\frac{1}{2\gamma}\|\theta^{k_{j}}-y^{k_{j}}\|^{2}\rightarrow v_{\gamma}(\bar{x},\bar{y})$ as $k_{j}\rightarrow\infty$ from the continuity of $\varphi$ and $\theta^{*}(x,y)$ (Proposition 2). Taking the limit as $k=k_{j}\rightarrow\infty$ in (41), we have
$\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})\leq\epsilon.$ (47)
Therefore, letting $\bar{\epsilon}:=\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})$, we have $\bar{\epsilon}\in[0,\epsilon]$ and together with (45), we obtain
[…]
which implies that $(\bar{x},\bar{y})$ is a KKT point of problem $({\rm VP})_{\gamma}^{\bar{\epsilon}}$ with $\bar{\epsilon}\leq\epsilon$. ∎
As established in Lemma 6, the Ostrowski condition $\lim_{k\to\infty}\|(x^{k+1},y^{k+1})-(x^{k},y^{k})\|=0$ is satisfied. By combining this result with Theorem 2 and [55, Proposition 8.3.10], we derive the following convergence result:
Corollary 3.
Let $\epsilon>0$. Suppose the sequence ${p_{k}}$ is bounded, and the sequence ${(x^{k},y^{k})}$ has an isolated accumulation point $(\bar{x},\bar{y})$, meaning there exists a neighborhood $Z$ of $(\bar{x},\bar{y})$ such that $(\bar{x},\bar{y})$ is the only accumulation point of ${(x^{k},y^{k})}$ in $Z$. Then the sequence ${(x^{k},y^{k})}$ converges to $(\bar{x},\bar{y})$, which is a KKT point of $({\rm VP})_{\gamma}^{\bar{\epsilon}}$ for some $\bar{\epsilon}\leq\epsilon$.
As demonstrated in Proposition 3, ENNAMCQ holds for $({\rm VP})_{\gamma}^{\epsilon}$ for any $\epsilon>0$ when the lower level problem is strictly convex with respect to the lower level variable $y$. Next, we show that ENNAMCQ is sufficient to guarantee the boundedness of the penalty parameter $p_{k}$ generated by Algorithm 1, as required in Theorem 2.
Theorem 4.
Let $\epsilon>0$. If the sequence $\left\{(x^{k},y^{k})\right\}$ is bounded, and ENNAMCQ holds at any accumulation points of $\left\{x^{k},y^{k}\right\}$, then $\left\{p_{k}\right\}$ must be bounded.
Proof.
The proof builds upon techniques used in the proof of Proposition 3 in [27]. We proceed by contradiction, assuming that $p_{k}\to\infty$ as $k\to\infty$. First, recall from Lemma 6 that $\lim_{k\to\infty}\|(x^{k+1},y^{k+1})-(x^{k},y^{k})\|=0$. From the update rule of $p_{k}$, there exist a subsequence $\{k_{j}\}\subseteq\mathbb{N}$ such that
$t^{k_{j}+1}>c_{p}^{-1}\left\|\left(x^{k_{j}},y^{k_{j}}\right)-\left(x^{k_{j}+1},y^{k_{j}+1}\right)\right\|>0.$ (49)
Since the sequence $\{(x^{k},y^{k})\}$ is bounded, we can assume without loss of generality by extracting a further subsequence that $(x^{k_{j}},y^{k_{j}})\to(\bar{x},\bar{y})$ for some $(\bar{x},\bar{y})\in X\times Y$. Suppose $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})-\epsilon<0$. As shown in the proof of Theorem 2, $\varphi(x^{k_{j}},\theta^{k_{j}})+\frac{1}{2\gamma}\|\theta^{k_{j}}-y^{k_{j}}\|^{2}\rightarrow v_{\gamma}(\bar{x},\bar{y})$ as $k_{j}\rightarrow\infty$. Thus, there exists $k_{j}$ such that
$\varphi(x^{k_{j}},y^{k_{j}})-\varphi(x^{k_{j}},\theta^{k_{j}})-\frac{1}{2\gamma}\left\|\theta^{k_{j}}-y^{k_{j}}\right\|^{2}-\epsilon<0,$
which implies $t^{k_{j}}=0$. This contradicts (49). Therefore, we conclude that $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})-\epsilon\geq 0$. Next, consider(42), which states
$p_{k}^{-1}\left(\xi_{x}^{k}-\nabla F(x^{k+1},y^{k+1})\right)$ $\in\nabla_{x}\varphi(x^{k+1},y^{k+1})-\nabla_{x}v_{\gamma}(x^{k+1},y^{k+1})+\mathcal{N}_{X}(x^{k+1}),$ (50)
$p_{k}^{-1}\left(\xi_{y}^{k}-\nabla F(x^{k},y^{k+1})\right)$ $\in\partial_{y}\varphi_{y}(x^{k},y^{k+1})-\nabla_{y}v_{\gamma}(x^{k},y^{k+1})+\mathcal{N}_{Y}(y^{k+1})$
where $\xi_{x}^{k},\ \xi_{y}^{k}$ are defined in (43). Using similar arguments as in the proof of Theorem 2, we can show that
$\lim_{k_{j}\to\infty}\left(p_{k_{j}}^{-1}\left(\xi_{x}^{k_{j}}-\nabla_{x}F(x^{{k_{j}}+1},y^{{k_{j}}+1})\right),p_{k_{j}}^{-1}\left(\xi_{y}^{k_{j}}-\nabla_{y}F(x^{{k_{j}}},y^{{k_{j}}+1})\right)\right)=0.$
Taking the limit as $k=k_{j}\rightarrow\infty$ in (50) yields
$0\in\nabla f(\bar{x},\bar{y})+\partial g(\bar{x},\bar{y})-\nabla v_{\gamma}(\bar{x},\bar{y})+\mathcal{N}_{X\times Y}(\bar{x},\bar{y}),$
which contradicts the assumption that ENNAMCQ holds and thus we get the conclusion. ∎
5 Sequential convergence under KL property
This section establishes the sequential convergence of AGILS (Algorithm 1) under the Kurdyka-Łojasiewicz (KL) property. We begin by recalling the definition of the KL property. Let $\zeta\in[0,+\infty)$ and $\Phi_{\zeta}$ represent the set of all concave and continuous functions $\phi\ :\ [0,\zeta)\to[0,+\infty)$ that meet the following conditions: (a) $\phi(0)=0$, (b) $\phi$ is $C^{1}$ on $(0,\zeta)$ and continuous at 0, (c) $\phi^{\prime}(s)>0$ for all $s\in(0,\zeta)$.
Definition 3 (Kurdyka-Łojasiewicz property).
Consider $h:{\mathbb{R}}^{d}\to(-\infty,+\infty]$ as proper and lower semicontinuous. The function $h$ is said to have the Kurdyka-Łojasiewicz (KL) property at $\bar{x}\in{\rm dom}\,\partial h:=\left\{x\in{\mathbb{R}}^{d}\ |\ \partial h(x)\neq\emptyset\right\}$, if there exist $\zeta\in(0,+\infty]$, a neighborhood $\mathcal{U}$ of $\bar{x}$ and a function $\phi\in\Phi_{\zeta}$, such that for all $x\in\mathcal{U}\bigcap\left\{x\in{\mathbb{R}}^{d}\ |\ h(\bar{x})<h(x)<h(\bar{x})+\zeta\right\}$, the following inequality holds
$\phi^{\prime}(h(x)-h(\bar{x})){\rm dist}(0,\partial h(x))\geq 1.$
If $h$ satisfies the KL property at each point of ${\rm dom}\,\partial h$ then $h$ is referred to as a KL function.
In addition, when the KL property holds for all points in a compact set, the uniformized KL property is applicable, refer to Lemma 6 in [52] for further details.
Lemma 7 (Uniformized KL property).
Given a compact set $D$ and a proper and lower semicontinuous function $h\ :\ {\mathbb{R}}^{d}\to(-\infty,+\infty]$, suppose that $h$ is constant on $D$ and satisfies the KL property at each point of $D$. Then, there exist $\epsilon$, $\zeta$ and $\phi\in\Phi_{\zeta}$ such that for all $\bar{x}\in D$ and $x\in\left\{x\in{\mathbb{R}}^{d}\,|\,{\rm dist}(x,D)<\epsilon,\ h(\bar{x})<h(x)<h(\bar{x})+\zeta\right\}$,
$\phi^{\prime}(h(x)-h(\bar{x})){\rm dist}(0,\partial h(x))\geq 1.$
To establish the sequential convergence of AGILS (Algorithm 1), we further assume that $\partial_{y}g(x,y)$ admits a Lipschitz continuity with respect to $x$ on $X$.
Assumption 4.
$\partial_{y}g(x,y)$ is $L_{g_{y}}$-Lipschitz continuous with respect to $x$ on $X$ for any fixed $y\in Y$. That is, for any $x,x^{\prime}\in X$,
$\partial_{y}g(x^{\prime},y)\subseteq\partial_{y}g(x,y)+L_{g_{y}}\|x-x^{\prime}\|\mathbb{B}.$
Assumption 4 is not restrictive. For example, if $g(x,y)=\sum_{i=1}^{p}Q_{i}(x)P_{i}(y)$, where $Q_{i}(x):X\rightarrow\mathbb{R}$ is Lipschitz continuous and strictly positive on $X$, and $P_{i}(y):Y\rightarrow\mathbb{R}$, then $g(x,y)$ satisfies Assumption 4. Specific examples include $g(x,y)=x\|y\|_{1}$ on ++×$\mathbb{R}_{++}\times\mathbb{R}^{m}$ and $g(x,y)=\sum_{j=1}^{J}x_{j}\|y^{(j)}\|_{2}$ on ++$\mathbb{R}^{J}_{++}\times\mathbb{R}^{m}$.
Consider the sequences $\{(x^{k},y^{k})\}$ and $\{p_{k}\}$ generated by AGILS (Algorithm 1) under the step size conditions specified in Section 4.2, where $\gamma\in(0,1/(\rho_{f_{2}}+\rho_{g_{2}}))$, the step sizes satisfy $\alpha_{k}\in[\underline{\alpha},2/(L_{\psi_{x,k}}+c_{\alpha})]$ and $\beta_{k}\in[\underline{\beta},2/(L_{\psi_{y,k}}+c_{\beta})]$, where $L_{\psi_{x,k}}=L_{F_{x}}/p_{k}+L_{f_{x}}+L_{g_{1}}+\rho_{f_{1}}+\rho_{g_{1}}$, $L_{\psi_{y,k}}=L_{F_{y}}/p_{k}+L_{f_{y}}$ and $\underline{\alpha},\underline{\beta},c_{\alpha},c_{\beta}$ are positive constants. To analyze sequential convergence, we define the following merit function $E_{p}(z,u)$, where $z:=(x,y)$,
$E_{p}(z,u):=G_{p}(z)-\langle u,x\rangle+H(u,y)$ (51)
with
$G_{p}(z):=\frac{1}{p}F(z)+\varphi(z)+\frac{\rho_{v_{1}}}{2}\|x\|^{2},\quad H(u,y):=\sup_{x}\{\langle u,x\rangle-\tilde{v}_{\gamma}(x,y)\},$
where $\rho_{v_{1}}=\rho_{f_{1}}+\rho_{g_{1}}+c_{\alpha}/4$ and $\tilde{v}_{\gamma}(x,y):={v}_{\gamma}(x,y)+(\rho_{v_{1}}/2)\|x\|^{2}$. By Proposition 2 and under Assumption 3, $\tilde{v}_{\gamma}(x,y)$ is strongly convex with respect to $x$. $H(x,u)$ is the Fenchel conjugate of $\tilde{v}_{\gamma}(x,y)$ with respect to variable $x$ and is differentiable due to the differentiability of ${v}_{\gamma}(x,y)$ and the strong convexity of ${v}_{\gamma}(x,y)$ with respect to $x$. For the sequence $\{(x^{k},y^{k})\}$ generated by AGILS(1), let $u^{k}=\nabla_{x}\tilde{v}_{\gamma}(x^{k},y^{k+1})$. We have $\nabla_{u}H(u^{k},y^{k+1})=x^{k}$ and $\nabla_{y}H(u^{k},y^{k+1})=\nabla_{y}\tilde{v}_{\gamma}(x^{k},y^{k+1})$. The following lemma establishes a relaxed sufficient decrease property and relative error condition for the merit function $E_{p}(z,u)$ at $(z^{k+1},u^{k})$.
Lemma 8.
Under Assumptions 1, 2, 3 and 4. Let $\epsilon>0$. Let $\{(x^{k},y^{k})\}$ and $\{p_{k}\}$ be the sequences generated by AGILS under the same conditions as in Section 4.2, and let $E_{p}(z,u)$ be defined as in (51), then there exist $a,b>0$ such that
$E_{p_{k}}(z^{k+1},u^{k})+a\left\|z^{k+1}-z^{k}\right\|^{2}\leq E_{p_{k}}(z^{k},u^{k-1})+\nu_{k},$ (52)
where $\nu_{k}=\left(\frac{1}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}+\frac{1}{c_{\beta}\gamma^{2}}\right)C_{\eta}^{2}s_{k}^{2}$, and
${\rm dist}\left(0,\partial E_{p_{k}}\left(z^{k+1},u^{k}\right)\right)\leq b\|z^{k+1}-z^{k}\|+\tilde{\nu}_{k},$ (53)
where $\tilde{\nu}_{k}=(L_{f_{x}}+L_{g_{2}}+1/\gamma)C_{\eta}s_{k}$.
Proof.
We begin by deriving the first inequality. Using (31), it follows that:
$G_{p_{k}}(x^{k},y^{k+1})-{v}_{\gamma}(x^{k},y^{k+1})$ (54)
$\leq$ $G_{p_{k}}(x^{k},y^{k})-{v}_{\gamma}(x^{k},y^{k})-\left(\frac{1}{\beta_{k}}-\frac{L_{\psi_{y,k}}}{2}-\frac{c_{\beta}}{4}\right)\left\|y^{k+1}-y^{k}\right\|^{2}+\frac{C_{\eta}^{2}}{c_{\beta}\gamma^{2}}s_{k}^{2}.$
Next, using (32) and the fact that the gradients $\nabla_{x}F(x,y^{k+1})$, $\nabla_{x}f(x,y^{k+1})$ and $\nabla_{x}g(x,y^{k+1})$ are $L_{F_{x}}$, $L_{f_{x}}$ and $L_{g_{1}}$-Lipschitz continuous with respect to variable $x$ on $X$, respectively, we have
$G_{p_{k}}(x^{k+1},y^{k+1})+\left(\frac{1}{\alpha_{k}}-\frac{L_{F_{x}}/p_{k}+L_{f_{x}}+L_{g_{1}}+\rho_{v_{1}}}{2}\right)\left\|x^{k+1}-x^{k}\right\|^{2}$ (55)
$\leq$ $G_{p_{k}}(x^{k},y^{k+1})+\left\langle\nabla_{x}G_{p_{k}}(x^{k},y^{k+1})-d_{x}^{k},x^{k+1}-x^{k}\right\rangle.$
Using the definitions of $G_{p}(z)$ and $\tilde{v}_{\gamma}(z)$, along with the gradient formulas for $\nabla_{x}{v}_{\gamma}(z)$ and $d_{x}^{k}$ in (19), we further derive
$G_{p_{k}}(z^{k+1})-\langle\nabla_{x}\tilde{v}_{\gamma}(x^{k},y^{k+1}),x^{k+1}-x^{k}\rangle+\left(\frac{1}{\alpha_{k}}-\frac{L_{E_{x,k}}}{2}\right)\left\|x^{k+1}-x^{k}\right\|^{2}$ (56)
$\leq$ $G_{p_{k}}(x^{k},y^{k+1})+\left\langle\nabla_{x}\varphi(x^{k},\theta^{k+1/2})-\nabla_{x}v_{\gamma}(x^{k},y^{k+1}),x^{k+1}-x^{k}\right\rangle,$
where $L_{E_{x,k}}=L_{F_{x}}/p_{k}+L_{f_{x}}+L_{g_{1}}+\rho_{v_{1}}$. Combining (35) with (56) and applying the Cauchy-Schwarz inequality, we obtain
$G_{p_{k}}(z^{k+1})-\langle\nabla_{x}\tilde{v}_{\gamma}(x^{k},y^{k+1}),x^{k+1}-x^{k}\rangle+\left(\frac{1}{\alpha_{k}}-\frac{L_{E_{x,k}}}{2}-\frac{c_{\alpha}}{4}\right)\left\|x^{k+1}-x^{k}\right\|^{2}$ (57)
$\leq$ $G_{p_{k}}(x^{k},y^{k+1})+\frac{1}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}C_{\eta}^{2}s_{k}^{2}.$
Combining this result with (54), we have
$G_{p_{k}}(z^{k+1})-\tilde{v}_{\gamma}(x^{k},y^{k+1})-\langle\nabla_{x}\tilde{v}_{\gamma}(x^{k},y^{k+1}),x^{k+1}-x^{k}\rangle$ (58)
$\leq$ $G_{p_{k}}(z^{k})-\tilde{v}_{\gamma}(z^{k})+\left(\frac{1}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}+\frac{1}{c_{\beta}\gamma^{2}}\right)C_{\eta}^{2}s_{k}^{2}$
$-\left(\frac{1}{\alpha_{k}}-\frac{L_{E_{x,k}}}{2}-\frac{c_{\alpha}}{4}\right)\left\|x^{k+1}-x^{k}\right\|^{2}-\left(\frac{1}{\beta_{k}}-\frac{L_{\psi_{y,k}}}{2}-\frac{c_{\beta}}{4}\right)\left\|y^{k+1}-y^{k}\right\|^{2}.$
We have $-\tilde{v}_{\gamma}(x^{k},y^{k+1})+\langle\nabla_{x}\tilde{v}_{\gamma}(x^{k},y^{k+1}),x^{k}\rangle=H(u^{k},y^{k+1})$, because $\tilde{v}_{\gamma}$ is convex in $x$ and $u^{k}=\nabla_{x}\tilde{v}_{\gamma}(x^{k},y^{k+1})$. Additionally, since $-\tilde{v}_{\gamma}(x^{k},y^{k})\leq-\langle u^{k-1},x^{k}\rangle+H(u^{k-1},y^{k})$, substituting this into (58) yields
$E_{p_{k}}(z^{k+1},u^{k})+c_{\alpha}/8\cdot\left\|x^{k+1}-x^{k}\right\|^{2}+c_{\beta}/4\cdot\left\|y^{k+1}-y^{k}\right\|^{2}\leq E_{p_{k}}(z^{k},u^{k-1})+\nu_{k},$ (59)
where $\nu_{k}=\left(\frac{1}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}+\frac{1}{c_{\beta}\gamma^{2}}\right)C_{\eta}^{2}s_{k}^{2}$, and (52) follows.
We now establish the relative error condition of $E_{p}(z,u)$. To begin, we provide the subdifferential characterization of $E_{p}(z,u)$,
$\partial E_{p}(z,u)=\begin{pmatrix}\nabla_{x}G_{p}(z)-u+\mathcal{N}_{X}(x)\\ \partial_{y}G_{p}(z)+\nabla_{y}H(u,y)+\mathcal{N}_{Y}(y)\\ \nabla_{u}H(u,y)-x\end{pmatrix}$ (60)
Using the update rules of $x^{k+1}$ and $y^{k+1}$ in (20) and (15), along with the formulas for $d_{x}^{k}$ and $d^{k}_{y}$ in (19) and (14), we have
00 0 $\in\frac{1}{p_{k}}\nabla_{x}F(x^{k},y^{k+1})+\nabla_{x}\varphi(x^{k},y^{k+1})-\nabla_{x}\varphi(x^{k},\theta^{k+1/2})+\frac{x^{k+1}-x^{k}}{\alpha_{k}}+\mathcal{N}_{X}(x^{k+1}),$
00 0 $\in\frac{1}{p_{k}}\nabla_{y}F(z^{k})+\nabla_{y}f(z^{k})+\partial_{y}g(x^{k},y^{k+1})-\frac{y^{k}-\theta^{k}}{\gamma}+\frac{y^{k+1}-y^{k}}{\beta_{k}}+\mathcal{N}_{Y}(y^{k+1}).$
From this, we have $\tilde{\xi}_{x}^{k}\in\nabla_{x}G_{p_{k}}(z^{k+1})-u^{k}+\mathcal{N}_{X}(x^{k+1})$ with
$\tilde{\xi}_{x}^{k}:=$ $\frac{1}{p_{k}}\left(\nabla_{x}F(z^{k+1})-\nabla_{x}F(x^{k},y^{k+1})\right)+\nabla_{x}\varphi(z^{k+1})-\nabla_{x}\varphi(x^{k},y^{k+1})$ (61)
$-\nabla_{x}v_{\gamma}(x^{k},y^{k+1})+\nabla_{x}\varphi(x^{k},\theta^{k+1/2})-\left(1/\alpha_{k}-\rho_{v_{1}}\right)(x^{k+1}-x^{k}).$
By Assumption 4, for any $\xi_{g}^{k}\in\partial_{y}g(x^{k},y^{k+1})$, there exists $\tilde{\xi}_{g}^{k}\in\mathbb{R}^{m}$ satisfying $\|\tilde{\xi}_{g}^{k}\|\leq L_{g_{y}}\|x^{k+1}-x^{k}\|$ such that ${\xi}_{g}^{k}+\tilde{\xi}_{g}^{k}\in\partial_{y}g(x^{k+1},y^{k+1})$. Consequently, we have $\tilde{\xi}_{y}^{k}\in\partial_{y}G_{p_{k}}(z^{k+1})+\nabla_{y}H(u^{k},y^{k+1})+\mathcal{N}_{Y}(y^{k+1})$, where
$\tilde{\xi}_{y}^{k}:=$ $\frac{1}{p_{k}}\left(\nabla_{y}F(z^{k+1})-\nabla_{y}F(z^{k})\right)+\nabla_{y}f(z^{k+1})-\nabla_{y}f(z^{k})+\tilde{\xi}_{g}^{k}$ (62)
$-\nabla_{y}v_{\gamma}(x^{k},y^{k+1})+(y^{k}-\theta^{k})/\gamma-(y^{k+1}-y^{k})/\beta_{k}.$
By the $L_{F_{x}}$-, $L_{f_{x}}$- and $L_{g_{1}}$-Lipschitz continuity of $\nabla_{x}F(x,y^{k+1})$, $\nabla_{x}f(x,y^{k+1})$ and $\nabla_{x}g(x,y^{k+1})$ with respect to variable $x$ on $X$, along with $\alpha_{k}\geq\underline{\alpha}$, we have
$\|\tilde{\xi}_{x}^{k}\|\leq$ $\left(L_{F_{x}}/p_{k}+L_{f_{x}}+L_{g_{1}}+1/\underline{\alpha}+\rho_{v_{1}}\right)\|x^{k+1}-x^{k}\|$ (63)
$+\|\nabla_{x}v_{\gamma}(x^{k},y^{k+1})-\nabla_{x}\varphi(x^{k},\theta^{k+1/2})\|$
$\leq$ $\left(L_{F_{x}}/p_{k}+L_{f_{x}}+L_{g_{1}}+1/\underline{\alpha}+\rho_{v_{1}}\right)\|x^{k+1}-x^{k}\|+(L_{f_{x}}+L_{g_{2}})C_{\eta}s_{k}$
where the second inequality follows from (35). For $\tilde{\xi}_{y}^{k}$, as $\nabla_{y}F(x^{k},y)$ and $\nabla_{y}f(x^{k},y)$ are $L_{F_{y}}$- and $L_{f_{y}}$-Lipschitz continuous on $X\times Y$, and $\beta_{k}\geq\underline{\beta}$, we have
$\|\tilde{\xi}_{y}^{k}\|\leq$ $\left(L_{F_{y}}/p_{k}+L_{f_{y}}+1/\underline{\beta}\right)\|z^{k+1}-z^{k}\|+\|\tilde{\xi}^{k}_{g}\|+\|\nabla_{y}v_{\gamma}(x^{k},y^{k+1})-(y^{k}-\theta_{y}^{k})/\gamma\|$ (64)
$\leq$ $\left(L_{F_{y}}/p_{k}+L_{f_{y}}+1/\underline{\beta}\right)\|z^{k+1}-z^{k}\|+L_{g_{y}}\|x^{k+1}-x^{k}\|$
$+\|\nabla_{y}v_{\gamma}(x^{k},y^{k+1})-\nabla_{y}v_{\gamma}(x^{k},y^{k})\|+\|\nabla_{y}v_{\gamma}(x^{k},y^{k})-(y^{k}-\theta_{y}^{k})/\gamma\|$
$\leq$ $\left(L_{F_{y}}/p_{k}+L_{f_{y}}+L_{g_{y}}+L_{\theta^{*}}/\gamma+1/\underline{\beta}+1/\gamma\right)\|z^{k+1}-z^{k}\|+{C_{\eta}}s_{k}/\gamma,$
where the last inequality follows from (30) and the $L_{\theta^{*}}$-Lipschitz continuity of $\theta^{*}(x,y)$ with respect to $y$ (from Lemma 2). For the last component of $\partial E_{p_{k}}(z^{k+1},u^{k})$, note that since $u^{k}=\nabla_{x}\tilde{v}_{\gamma}(x^{k},y^{k+1})$, we have $\nabla_{u}H(u^{k},y^{k+1})=x^{k}$. Thus
$\|\nabla_{u}H(u^{k},y^{k+1})-x^{k+1}\|=\|x^{k}-x^{k+1}\|.$ (65)
Finally, combining (63), and (64), and using the fact that $(\tilde{\xi}_{x}^{k},\tilde{\xi}_{y}^{k},x^{k}-x^{k+1})\in\partial E_{p_{k}}(z^{k+1},u^{k})$, the desired result (53) follows. ∎
To establish sequential convergence under the relaxed conditions (52) and (53), we impose a stronger requirement on the inexact parameter $s_{k}$. Specifically, $s_{k}$ satisfies
$\sum_{k=1}^{\infty}s_{k}<\infty\quad\text{and}\quad\sum_{k=1}^{\infty}ks_{k}^{p_{s}}<\infty,$ (66)
for some $p_{s}<2$. For example, $s_{k}$ can be chosen as $s_{k}=1/k^{1.1}$. Under this condition, there exists $q>0$ such that $2(1-1/q)>p_{s}$, ensuring that $\sum_{k=1}^{\infty}ks_{k}^{2(1-1/q)}<\infty$. To facilitate the convergence analysis, we introduce the following auxiliary function,
$\tilde{E}_{p}(z,u,w)=E_{p}(z,u)+w^{q},$
where $E_{p}(z,u)$ is as defined in (51). Since $w^{q}$ is a KL function, it follows from [56, Theorem 3.6] (although this result considers the generalized concave KL property, it can be directly extended to the KL property using the same proof) that if $E_{p}(z,u)$ is a KL function, then $\tilde{E}_{p}(z,u,w)$ is also a KL function.
Theorem 5.
Under Assumptions 1, 2, 3 and 4, let $\epsilon>0$. Consider the sequences $\{(x^{k},y^{k})\}$ and $\{p_{k}\}$ generated by AGILS (1) under the same conditions as in Section 4.2, with $s_{k}$ satisfying condition (66). Suppose the sequence $\left\{(x^{k},y^{k})\right\}$ is bounded and $p_{k}=\bar{p}$ for sufficiently large $k$ and $E_{\bar{p}}(z,u)$ is a KL function. Then the sequence $\left\{x^{k},y^{k}\right\}$ converges to a KKT point of $({\rm VP})_{\gamma}^{\tilde{\epsilon}}$ for some $\tilde{\epsilon}\leq\epsilon$.
Proof.
Since there exists $k_{0}>0$ such that $p_{k}=\bar{p}:=p_{k_{0}}$ for any $k\geq k_{0}$. Without loss of generality, we assume that $p_{k}=\bar{p}$ for all $k$. From Lemma 6, we know that $\lim_{k\rightarrow\infty}\|z^{k+1}-z^{k}\|^{2}=0$. Let $\Omega$ denote the set of all limit points of the sequence $\{(z^{k},u^{k-1})\}$. From Proposition 2, we know that $\nabla v_{\gamma}$ is continuous. Given that the sequence $\{z^{k}\}$ is bounded, and $u^{k}=\nabla_{x}v_{\gamma}(x^{k},y^{k+1})+\rho_{v_{1}}x^{k}$, we conclude that the sequence $\{u^{k}\}$ is also bounded. Consequently, $\Omega$ is a bounded set. Since limit points form a closed set, it follows that $\Omega$ is compact and $\lim_{k\rightarrow\infty}\mathrm{dist}((z^{k},u^{k-1}),\Omega)=0$. Now, define $w^{k}=(\sum_{i=k}^{\infty}\nu_{i})^{1/q}$, since $\sum_{k}\nu_{k}<\infty$, we have $w^{k}\rightarrow 0$. We can also have $\tilde{E}_{\bar{p}}(z^{k},u^{k-1},w^{k})=E_{\bar{p}}(z^{k},u^{k-1})+\sum_{i=k}^{\infty}\nu_{i}$. By Lemma 8, there exist constants $a,b>0$ such that
$\tilde{E}_{\bar{p}}(z^{k+1},u^{k},w^{k+1})+a\left\|z^{k+1}-z^{k}\right\|^{2}\leq\tilde{E}_{\bar{p}}(z^{k},u^{k-1},w^{k}),\quad\text{and}$ (67)
${\rm dist}\left(0,\tilde{E}_{\bar{p}}\left(z^{k+1},u^{k},w^{k+1}\right)\right)\leq b\|z^{k+1}-z^{k}\|+\tilde{\nu}_{k}+q|w^{k+1}|^{q-1}.$ (68)
Since $\tilde{E}_{\bar{p}}(z,u,w)$ is bounded below and continuous on $X\times Y$, $\lim_{k\rightarrow\infty}\|z^{k+1}-z^{k}\|=0$, and $\tilde{E}_{\bar{p}}(z^{k},u^{k-1},w^{k})$ is nonincreasing as $k$ increases, therefore, there exists a constant $\bar{E}$ such that for any subsequence $\{(z^{k_{j}},u^{k_{j}-1},w^{k_{j}})\}$ of sequence $\{(z^{k},u^{k-1},w^{k})\}$, we have $\bar{E}=\lim_{k_{j}\rightarrow\infty}\tilde{E}_{\bar{p}}(z^{k_{j}},$ $u^{k_{j}-1},w^{k_{j}})=\lim_{k\rightarrow\infty}\tilde{E}_{\bar{p}}(z^{k},u^{k-1},w^{k})$. This implies that the function $\tilde{E}_{\bar{p}}(z,u,w)$ is constant on $\Omega\times\{0\}$. We can assume that $\tilde{E}_{\bar{p}}(z^{k},u^{k-1},w^{k})>\bar{E}$ for all $k$. If there exists some $k>0$ such that $\tilde{E}_{\bar{p}}(z^{k},u^{k-1},w^{k})=\bar{E}$, then from (67), we would have $\|z^{k+1}-z^{k}\|=0$ for sufficiently large $k$, implying that the sequence $\{z^{k}\}$ converges and the conclusion follows immediately.
Since $\lim_{k\rightarrow\infty}\mathrm{dist}((z^{k},u^{k-1},w^{k}),\Omega\times\{0\})=0$ and $\lim_{k\rightarrow\infty}\tilde{E}_{\bar{p}}(z^{k},u^{k-1},w^{k})=\bar{E}$, for any $\epsilon,\eta>0$, there exists $k_{0}$ such that $\mathrm{dist}((z^{k},u^{k-1},w^{k}),\Omega\times\{0\})<\epsilon$ and $\bar{E}<\tilde{E}_{\bar{p}}(z^{k},u^{k-1},w^{k})<\bar{E}+\eta$ for all $k\geq k_{0}$. Since $\tilde{E}_{\bar{p}}(z,u,w)$ satisfies the Kurdyka-Łojasiewicz property at each point in $\Omega\times\{0\}$, and $\tilde{E}_{\bar{p}}(z,u,w)$ is equal to a finite constant on $\Omega\times\{0\}$, we can apply Lemma 7 to obtain a continuous concave function $\phi$ such that for all $k\geq k_{0}$,
$\phi^{\prime}\left(\tilde{E}_{\bar{p}}(z^{k},u^{k-1},w^{k})-\bar{E}\right)\mathrm{dist}\left(0,\partial\tilde{E}_{\bar{p}}(z^{k},u^{k-1},w^{k})\right)\geq 1.$
For notational simplicity, we define $\tilde{E}_{k}:=\tilde{E}_{\bar{p}}(z^{k},u^{k-1},w^{k})-\bar{E}$. By combining the above inequality with (68), we obtain the following inequality,
$\phi^{\prime}\left(\tilde{E}_{k}\right)\cdot\,\left(b\|z^{k}-z^{k-1}\|+\tilde{\nu}_{k-1}+q|w^{k}|^{q-1}\right)\geq 1.$
Next, using the concavity of $\phi$ and (67), we obtain
$\phi^{\prime}(\tilde{E}_{k})\cdot a\|z^{k+1}-z^{k}\|^{2}\leq\phi^{\prime}(\tilde{E}_{k})\cdot\left(\tilde{E}_{k}-\tilde{E}_{k+1}\right)\leq\phi\left(\tilde{E}_{k}\right)-\phi\left(\tilde{E}_{k+1}\right).$
Combining the above two inequalities, we get
$\left(b\|z^{k}-z^{k-1}\|+\tilde{\nu}_{k-1}+q|w^{k}|^{q-1}\right)\left[\phi\left(\tilde{E}_{k}\right)-\phi\left(\tilde{E}_{k+1}\right)\right]$
$\geq$ $\left(b\|z^{k}-z^{k-1}\|+\tilde{\nu}_{k-1}+q|w^{k}|^{q-1}\right)\cdot\phi^{\prime}(\tilde{E}_{k})\cdot a\|z^{k+1}-z^{k}\|^{2}\geq a\|z^{k+1}-z^{k}\|^{2}.$
Multiplying both sides of this inequality by $4/a$ and then taking the square root, we apply the inequality $2cd\leq c^{2}+d^{2}$ to obtain
$2\|z^{k+1}-z^{k}\|\leq\|z^{k}-z^{k-1}\|+\tilde{\nu}_{k-1}/b+q|w^{k}|^{q-1}/b+b/a\cdot\left[\phi\left(\tilde{E}_{k}\right)-\phi\left(\tilde{E}_{k+1}\right)\right].$
Summing up the above inequality for $k=k_{0},\ldots,K$, since $\phi\geq 0$, we have
$\sum_{k=k_{0}}^{K}\|z^{k+1}-z^{k}\|\leq\|z^{1}-z^{0}\|+\frac{b}{a}\phi\left(\tilde{E}_{1}\right)+\sum_{k=k_{0}}^{K}\frac{\tilde{\nu}_{k-1}}{b}+\sum_{k=k_{0}}^{K}\frac{q|w^{k}|^{q-1}}{b}.$
Taking $K\rightarrow\infty$ in the above inequality, and using the fact that $\sum_{k=1}^{\infty}\tilde{\nu}_{k-1}<\infty$ and $\sum_{k=1}^{\infty}|w^{k}|^{q-1}=\sum_{k=1}^{\infty}|(\sum_{i=k}^{\infty}\nu_{i})^{1/q}|^{q-1}\leq\sum_{k=1}^{\infty}\sum_{i=k}^{\infty}\nu_{i}^{1-1/q}=\sum_{k=1}^{\infty}k\nu_{i}^{1-1/q}\leq C_{\nu}(\sum_{k=1}^{\infty}ks_{k}^{2(1-1/q)})<\infty$ for some $C_{\nu}>0$, we obtain
$\sum_{k=1}^{\infty}\|z^{k+1}-z^{k}\|<\infty.$
Thus, the sequence $\{z^{k}\}$ is a Cauchy sequence. Hence the sequence $\{z^{k}\}$ converges and we get the conclusion from Theorem 2
Remark 2.
To further substantiate the applicability of our sequential convergence theory in applications, we now discuss the KL property imposed on the merit function $E_{\bar{p}}(z,u)$ in Theorem 5. A wide range of functions automatically satisfy the KL property, see, e.g., [57, 58, 59, 52]. Notably, as shown in [60, 61], semi-algebraic function is an important subclass of KL functions. More importantly, the class of semi-algebraic functions is closed under various operations, see, e.g., [57, 58, 52]. Specifically, the indicator functions of semi-algebraic sets, finite sum and product of semi-algebraic functions, composition of semi-algebraic functions and partial minimization of semi-algebraic function over semi-algebraic set are all semi-algebraic functions.
Based on these properties, if the objective functions $f$ and $g$ and the constraint set $Y$ in the lower-level problem are semi-algebraic, the Moreau envelope function $v_{\gamma}$ is also semi-algebraic. Additionally, if $F$ and the constraint set $X$ in the upper-level are semi-algebraic, the merit function $E_{\bar{p}}(z,u)$ defined in (51) is semi-algebraic and thus satisfies the KL property. Therefore, when $F$, $f$, $g$, $X$ and $Y$ in (1) are all semi-algebraic, the KL property of $E_{\bar{p}}(z,u)$ holds, ensuring the sequential convergence described in Theorem 5.
6 Numerical experiments
In this section, we evaluate the performance of our proposed algorithm, AGILS, by comparing it with several existing methods on two problems: a toy example and the sparse group Lasso bilevel hyperparameter selection problem. Both problems are specific instances of bilevel problem (1) and satisfy Assumptions 1, 2, 3, and 4. The baseline methods for comparison include: grid search, random search, Tree-structured Parzen Estimator (TPE [62] [^1] https://github.com/hyperopt/hyperopt, a Bayesian optimization method), Implicit Gradient-based Joint Optimization (IGJO [32] [^2] https://github.com/jjfeng/nonsmooth-joint-opt, an implicit differentiation method), and Moreau Envelope-based Hessian-free Algorithm (MEHA [43], a single-loop, gradient-based method), Value Function-Based Difference-of-Convex Algorithm (VF-iDCA [28] [^3] https://github.com/SUSTech-Optimization/VF-iDCA, a method based on a difference-of-convex algorithm ), and the MPCC approach (which reformulates the bilevel problem as a mathematical program with complementarity constraints). Methods that require smoothness assumptions, such as those in [34, 39], were excluded because the tested problems do not satisfy the required smoothness conditions. For grid search, random search, and TPE, each subproblem was solved using the CVXPY package with the CLARABEL and SCS solvers. All experiments were implemented in Python and conducted on a laptop with an Intel i7-1260P CPU (2.10 GHz) and 32 GB RAM.
6.1 Toy example
We consider the following toy example:
$\min\limits_{x\in{\mathbb{R}}^{n},y\in{\mathbb{R}}^{n},0\leq x\leq 1}\sum\limits_{i=0}^{n}y_{i}\quad{\rm s.t.}~{}~{}y\in\underset{y\in{\mathbb{R}}^{n}}{\mathrm{argmin}}\ \sum\limits_{i=0}^{n}\sqrt{(y_{i}-a_{i})^{2}+\frac{1}{n^{2}}}+\sum\limits_{i=0}^{n}x_{i}|y_{i}|,$ (69)
where $a\in{\mathbb{R}}^{n}$ satisfies $a_{i}=-2/n^{2/3}$ for $i\in 1,\dots,n/2$, and $a_{i}=-2/n^{2/3}$ for $i\in n/2+1,\ldots,n$. The optimal solution is $x^{*}_{i}=0$, $y^{*}_{i}=a_{i}$ for $i=1,\dots,n/2$ and $x^{*}_{i}\in[a_{i}/\sqrt{a_{i}^{2}+1/n^{2}},1]$, $y^{*}_{i}=0$ for $i=n/2+1,\dots,n$.
We compare the performance of AGILS with several methods, including grid search, random search, TPE, MEHA, and the MPCC approach. The performance of each method is evaluated using metric Error $:={\rm dist}(z^{k},S^{*})/\sqrt{1+\min_{z\in S^{*}}\lVert z^{*}\rVert^{2}}$, where $S^{*}$ is the optimal set of (69), and $z:=(x,y)$. For grid search, we set $x_{1}=x_{2}=\dots=x_{n}=\rho$, and perform the search over $\rho$ on a grid of 100 uniformly spaced values within the interval $[0,1]$. Random search is implemented by uniformly sampling 100 points from $[0,1]^{n}$. For TPE, the search space is modeled as a uniform distribution over $[0,1]^{n}$. In the MPCC approach, we first reformulate the toy example by introducing non-negative vectors $u$ and $v$ to represent $y=u-v$, and transform $\|y\|_{1}$ into $\sum_{i=1}^{n}(u_{i}+v_{i})$. Then the corresponding MPCC reformulation is solved using the IPOPT solver with a relaxation parameter of 0.1 on the complementarity constraint and a solver tolerance of 0.1. For MEHA, the initial iterates are the same as AGILS, and the parameters are searched over $\tilde{c}_{0}\in\{0.1,0.5,1,5\}$, $p\in\{0.1,0.2,0.3,0.4\}$ and $\tilde{\alpha}$, $\tilde{\beta}$, $\tilde{\eta}$, and $\tilde{\gamma}$ scaled proportionally from the AGILS parameters ${\alpha}$, ${\beta}$, ${\eta}$, and ${\gamma}$ using scales $\{1,1/2,1/3,1/4\}$. The optimal parameters that minimize the Error are selected as $\tilde{c}_{0}=1$, $p=0.2$, $\tilde{\gamma}=\gamma/3$, $\tilde{\alpha}=\alpha/4$, $\tilde{\beta}=\beta$, and $\tilde{\eta}=\eta$. MEHA terminates when $\mathrm{Error}<1/n$. For AGILS, the initial iterates are $x_{0}=y_{0}=\theta_{0}=[0,\dots,0]$, the parameters are set as $\epsilon=10^{-6}$, $s_{k}=0.05/(k+1)^{1.05},$ and $p_{0}=0.5,\ \varrho_{p}=0.02,\ c_{p}=1$. The step sizes $\alpha,\beta,\gamma$ are chosen as $\alpha_{k}=1/(L_{\psi_{x,k}}+c_{\alpha})$, $\beta_{k}=1/(L_{\psi_{y,k}}+c_{\beta})$, and $\gamma=1/(\rho_{f_{2}}+\rho_{g_{2}})$ where $L_{\psi_{x,k}}=L_{F_{x}}/p_{k}+L_{f_{x}}+L_{g_{1}}+\rho_{f_{1}}+\rho_{g_{1}}$, $L_{\psi_{y,k}}=L_{F_{y}}/p_{k}+L_{f_{y}}$ with the following values for the constants $L_{F_{x}}=L_{F_{y}}=L_{f_{x}}=\rho_{f_{1}}=\rho_{f_{2}}=L_{g_{1}}=0$, $\rho_{g_{1}}=\rho_{g_{2}}=1$, $L_{f_{y}}=n$, $c_{\alpha}=c_{\beta}=0.1$ and $\eta=1/(L_{f_{y}}+1/\gamma)$. AGILS terminates when $\text{Error}<1/n$. The approximations $\theta^{k}$ are found at each iteration using the proximal gradient method to satisfy inexact criteria (18) and (21).
Numerical results for two different problem dimensions, $n=200$ and $n=600$, are reported in Table 1. The results demonstrate that AGILS performs well in both computational efficiency and solution accuracy, consistently achieving the lowest Error with the shortest computational time. MEHA performs second, but it requires tuning of parameters from a set of candidate values, which can be a limitation.
Table 1: Comparisons of different methods on the toy example for dimensions $n=200$ and $n=600$
Dimension $n=200$
Method Grid Random TPE MPCC MEHA AGILS
Time (s) 0.62 4.83 28.55 19.23 0.08 0.08
Error 0.70 0.75 0.72 0.02 0.00 0.00
Dimension $n=600$
Method Grid Random TPE MPCC MEHA AGILS
Time(s) 2.14 25.72 98.92 990.91 0.42 0.14
Error 0.71 0.79 0.76 0.27 0.00 0.00
Next, we evaluate the effectiveness of the proposed inexact criterion (18) and (21) by comparing AGILS with two extreme variants that use different inexact criteria on the toy example with $n=200$. One variant, denoted as AGILS-E, solves the proximal lower-level problem almost exactly, using a very small tolerance $10^{-6}$ in (18) and (21), while the other, denoted as AGILS-S, solves the proximal lower-level problem using only a single proximal gradient step. The parameters and tolerances used are consistent with those described above. The results in Figure 1 show that AGILS-E achieves the fastest error reduction in terms of iterations, but is slow in runtime due to the high computational cost of solving the proximal lower-level problem exactly. AGILS-S fails to converge due to its excessive inexactness in solving the proximal lower-level problem. AGILS achieves the fastest error reduction in terms of runtime.
Figure 1: Effectiveness of the inexact criterion in AGILS: comparison with two extreme variants
Finally, we evaluate the scalability of AGILS by testing it on toy examples with varying problem dimensions. The parameters of AGILS are as specified earlier, except that $p_{0}$ is set as 1. AGILS terminates when $\text{Error}<1/n$. The results in Figure 2 show that computational time increases steadily as problem dimension grows. This demonstrates that AGILS remains both efficient and stable as the problem dimension increases, highlighting its ability to handle large-scale problems effectively.
Figure 2: Iteration and computational time of AGILS on the toy example with varying dimensions
6.2 Sparse group Lasso
The sparse group Lasso [63] is a regularized regression model designed to achieve both individual feature sparsity and group-level sparsity. Given $m$ features divided into $J$ groups, let $y^{(j)}$ and $x_{j}$ denote the coefficients and regularization parameter of the $j$-th group. The corresponding bilevel hyperparameter selection problem is formulated as follows:
$\min_{x\in{{\mathbb{R}}}^{J+1}_{+},y\in{{\mathbb{R}}}^{m}}$ $\frac{1}{2|I_{\text{val}}|}\sum_{i\in I_{\text{val}}}\lvert b_{i}-y^{T}a_{i}\rvert^{2}$
${\rm s.t.}$ $y\in\underset{\hat{y}\in{{\mathbb{R}}}^{m}}{\mathrm{argmin}}\,\frac{1}{2|I_{\text{tr}}|}\sum_{i\in I_{\text{tr}}}\lvert b_{i}-\hat{y}^{T}a_{i}\rvert^{2}+\sum_{j=1}^{J}x_{j}\lVert\hat{y}^{\left(m\right)}\rVert_{2}+x_{J+1}\lVert\hat{y}\rVert_{1},$
where $I_{\text{val}}$ and $I_{\text{tr}}$ are the sets of validation and training sample indices, respectively. In this part, we compare the performance of our AGILS algorithm with several hyperparameter optimization methods, including grid search, random search, TPE, IGJO, VF-iDCA, and MEHA. The MPCC approach is excluded from the comparison due to its prohibitively high computational runtime. Following [32], the data generation process is as follows: $a_{i}\in\mathbb{R}^{m}$ is sampled from $N(0,I)$, and the response $b_{i}$ is computed as $b_{i}=y^{T}a_{i}+\sigma\epsilon_{i}$, where noise $\epsilon$ is sampled from $N(0,1)$. The coefficient vector $y$ is divided equally into five groups, with the $i$th group having the first $2i$ features set to value $2i$ and the remaining features set to zero, for $i=1,\ldots,5$. The noise level $\sigma$ is chosen to ensure $\text{SNR}=\|Ay\|/\|b-Ay\|=3$. The dataset is then randomly split into training, validation, and test sets of sizes $n_{\text{tr}}$, $n_{\text{val}}$, and $n_{\text{test}}$, respectively.
For grid search, random search and TPE, we parameterize $x_{i}$ with the log-transformed variables $\rho_{i}=\log_{10}(x_{i})$ and the search range is $[-9,2]^{J+1}$. For grid search, we set $\rho_{1}=\rho_{2}=\dots=\rho_{M}$ and use a $20\times 20$ grid. Random search samples 400 points uniformly within the range. For TPE, the search space is modeled as a uniform distribution over the same range. For VF-iDCA, MEHA, IGJO, and AGILS, the initial value of $x_{0}$ is set to $[1,\dots,1]$. Both $y_{0}$ and $\theta_{0}$ for MEHA and AGILS are initialized as $[1,\dots,1]$. For VF-iDCA, the parameters are selected as $\beta_{0}=5$, $\rho=0.01$, $c=1$, and $\delta=0.5$. VF-iDCA terminates when $\max\{\lVert{z}^{k+1}-{z}^{k}\rVert/\sqrt{1+\lVert{z}^{k}\rVert^{2}},\tilde{t}^{k+1}/m\}<0.1$, where $\tilde{t}^{k+1}$ is defined in [28]. The IGJO algorithm is run with a maximum of 50 iterations and a minimum step size of $10^{-6}$. For MEHA, the parameters are searched over $\tilde{c}_{0}\in\{1,10,20\}$ and $p\in\{0.1,0.2,0.3,0.4\}$, while $\tilde{\alpha}$, $\tilde{\beta}$, $\tilde{\eta}$, and $\tilde{\gamma}$ scaled proportionally from the AGILS parameters ${\alpha}$, ${\beta}$, ${\eta}$, and ${\gamma}$ using scales $\{1,1/2,1/4,1/8\}$. The optimal parameters that minimize the validation error are selected as $\tilde{c}_{0}=20$, $p=0.1$, $\tilde{\gamma}=\gamma$, $\tilde{\alpha}=\alpha$, $\tilde{\beta}=\beta/9$, and $\tilde{\eta}=\eta/5$. MEHA terminates when $\lVert\mathbf{z}^{k+1}-\mathbf{z}^{k}\rVert/\sqrt{1+\lVert\mathbf{z}^{k}\rVert^{2}}<0.005/m$. AGILS uses the same parameter settings as in the toy example, with specific values for constants such as $L_{F_{x}}=L_{f_{x}}=\rho_{f_{1}}=\rho_{f_{2}}=L_{g_{1}}=0$, $L_{F_{y}}=\lambda_{\text{max}}(A_{\text{val}}^{\top}A_{\text{val}})/n_{\text{val}}$, $L_{f_{y}}=\lambda_{\text{max}}(A_{\text{tr}}^{\top}A_{\text{tr}})/n_{\text{tr}}$, $\rho_{g_{1}}=1$, $\rho_{g_{2}}=m$. The penalty parameters are set as $p_{0}=6$, $\varrho_{p}=0.01$. AGILS terminates when $\lVert z^{k+1}-z^{k}\rVert/\sqrt{1+\lVert z^{k}\rVert^{2}}<0.005/m$, $z^{k}:=(x^{k},y^{k})$ and $t^{k+1}<0.1$. Proximal gradient method is used to compute $\theta^{k}$ that satisfies inexact criteria (18) and (21).
The experiments are repeated 20 times with $n_{\text{tr}}=200$, $n_{\text{val}}=200$, $n_{\text{test}}=200$, and $m=300$. Results are reported in Table 2 and Figure 3. We report the validation error (“Val. Err.”) and test error (“Test. Err.”), computed as $\sum_{i\in I_{\text{val}}}\lvert b_{i}-y^{T}a_{i}\rvert^{2}/|I_{\text{val}}|$ and $\sum_{i\in I_{\text{test}}}\lvert b_{i}-y^{T}a_{i}\rvert^{2}/|I_{\text{test}}|$, respectively. Both errors are evaluated using the sparse group Lasso estimator, which is obtained by solving the lower-level problem with the hyperparameter value corresponding to the iterates produced by the algorithms. Additionally, we report a second type of test error, referred to as the infeasible test error (“Test. Err. Infeas.”), where the lower-level variable values are taken directly from the iterates generated by VF-iDCA, MEHA and AGILS. “Feasibility” is a scaled measure of the value function constraint violation, given by $(\varphi(x^{k},y^{k})-v(x^{k}))/|I_{\text{val}}|$ for VF-iDCA, and $(\varphi(x^{k},y^{k})-v_{\gamma}(x^{k},y^{k}))/|I_{\text{val}}|$ for MEHA and AGILS.
Table 2: Comparison of different methods on the sparse group Lasso bilevel hyperparameter selection problem with $n_{\text{tr}}=n_{\text{val}}=n_{\text{test}}=200,m=300$
Method Time(s) Val. Err. Test. Err. Test. Err. Infeas. Feasibility
Grid 50.91(1.79) 168.26(23.91) 170.10(21.28) - -
Random 81.53(3.76) 185.87(27.24) 193.93(26.39) - -
TPE 137.73(4.45) 168.16(21.33) 178.96(22.70) - -
IGJO 192.95(99.66) 139.43(12.32) 169.62(21.11) - -
VF-iDCA 88.59(37.66) 134.68(14.71) 150.40(12.02) 138.70(11.77) 0.04(0.03)
MEHA 16.18(3.77) 98.35(8.79) 158.61(13.62) 158.42(13.80) 0.00(0.00)
AGILS 14.93(4.31) 96.06(10.53) 155.25(12.48) 155.39(12.05) 0.00(0.00)
Figure 3: Comparison of different methods on the sparse group Lasso bilevel hyperparameter selection problem with $n_{\text{tr}}=n_{\text{val}}=n_{\text{test}}=200,m=300$. Left: Val. Err., Right: Test. Err. Infeas.
As shown in Table 2, our proposed method, AGILS, outperforms the others in terms of validation error while requiring the least computation time. The validation error, corresponding to the value of the upper-level objective at a feasible point derived from the hyperparameter value of the iterates, can be seen as an indicator of optimization performance. AGILS also achieves competitive test errors. Interestingly, VF-iDCA finds the parameter $y$ that has the lowest test error, but the corresponding constraint violation is large. MEHA performs similarly to AGILS, but slightly less effective. Moreover, MEHA’s performance is highly sensitive to algorithmic parameters, and selecting inappropriate parameters can significantly degrade its effectiveness. Figure 3 illustrates that AGILS has fast convergence during the initial iterations.
Next, we investigate the impact of different solvers for solving the proximal lower-level problem in AGILS. We compare three algorithms: the Proximal Gradient Method (PGM), the Fast Iterative Shrinkage-Thresholding Algorithm (FISTA), and the Alternating Direction Method of Multipliers (ADMM). Each experiment is repeated 20 times, with all other parameters of AGILS held consistent with those used above. The results, presented in Table 3, show that the AGILS framework is robust to different solvers, with each solver exhibiting similar performance. This supports the flexibility and adaptability of the AGILS framework.
Table 3: Comparison of AGILS with different proximal lower-level problem solvers on the sparse group Lasso bilevel hyperparameter selection problem with $n_{\text{tr}}=n_{\text{val}}=n_{\text{test}}=200,m=300$
Method Time(s) Val. Err. Test. Err. Test. Err. Infeas. Feasibility
AGILS(PGM) 14.93(4.31) 96.06(10.53) 155.25(12.48) 155.39(12.05) 0.00(0.00)
AGILS(FISTA) 15.82(5.07) 96.06(10.53) 155.25(12.48) 155.39(12.05) 0.00(0.00)
AGILS(ADMM) 15.15(4.38) 95.60(10.11) 154.27(12.30) 154.58(12.11) 0.00(0.00)
Finally, we test the performance of AGILS on larger-scale problems with varying dimensions. The experiments, each repeated five times, are reported in Table 4. AGILS is terminated when $\lVert z^{k+1}-z^{k}\rVert/\sqrt{1+\lVert z^{k}\rVert^{2}}<0.1/m$ and $t^{k+1}<0.1$. All other parameters of AGILS remain consistent with the settings described above. The results demonstrate AGILS’s ability to efficiently solve larger-scale problems.
Table 4: AGILS on large-scale sparse group Lasso bilevel hyperparameter selection problems
Dimension Time(s) Val. Err. Test. Err. Test. Err. Infeas. Feasibility
$n_{\text{tr}}=1000,\ n_{\text{val}}=1000$  $\ n_{\text{test}}=1000,\ m=1500$ 20.13(3.72) 113.07(5.87) 116.85(9.54) 114.33(9.08) 0.00(0.00)
$n_{\text{tr}}=3000,\ n_{\text{val}}=3000$  $\ n_{\text{test}}=3000,\ m=4500$ 131.14(16.12) 106.91(2.92) 112.13(3.06) 109.95(2.70) 0.00(0.00)
$n_{\text{tr}}=5000,\ n_{\text{val}}=5000$  $\ n_{\text{test}}=5000,\ m=7500$ 433.96(35.61) 107.95(2.45) 106.64(1.50) 105.28(1.51) 0.00(0.00)
$n_{\text{tr}}=7000,\ n_{\text{val}}=7000$  $\ n_{\text{test}}=7000,\ m=10500$ 839.55(106.05) 105.71(1.80) 105.91(2.15) 104.72(2.14) 0.00(0.00)